### v1

# The Study of the Canonical forms of Killing tensor in vacuum with $\Lambda$

Abstract
This paper is the initial part of a comprehensive study of spacetimes that admit the canonical forms of Killing tensor in General Relativity. Our scope is to derive either new exact solutions of Einstein’s equations or to determine the hidden symmetries of the already known ones. In this preliminary work we first introduce the canonical forms of Killing tensor. Subsequently, we employ the integrability conditions of each canonical form along with the Einstein field equations (in vacuum with $\Lambda$) and the Bianchi identities in an attempt to create a solvable yet overdetermined system of equations. Finally, we obtain multiple special algebraic solutions according to the Petrov classification (D, III, N, O). The latter becomes possible since our analysis is embodied with the usage of the Newman-Penrose formalism of null tetrads.
Keywords: Canonical forms, Killing tensor, Newman-Penrose formalism, Exact solutions,
Einstein’s equations, Type D, Type III, Type N
1 Introduction
The interpretation and the analysis of exact solutions in the context of General Relativity constitutes a whole regime of research. It is already within our understanding that the Einstein’s Field Equations (EFE) cannot be solved without additional mathematical assumptions regarding the nature of a spacetime. Thus, the consideration of symmetries in the resolution process of the EFE is indispensable. In other words, the non-linear character of the equations obligate us to introduce further information through symmetries to obtain a solvable-overdetermined system. In this work we are not interested in any kind of symmetry, but we put under the spotlight spacetimes with hidden symmetries. We aim to find spacetimes which admit Killing Tensors since they are responsible for both explicit and hidden symmetries.
A well known example of a hidden symmetry is the conservation of Laplace-Runge-Lenz vector in the Kepler-Coulomb problem along geodesics as a constant of motion which is connected with the existence of Killing-Yano Tensor [Papakostas, 2001], a generalization of Killing Tensor [Cariglia, 2014],[Papakostas, 1985]. Also, a relevant conjecture says that: in a spacetime, which admit a non-trivial Killing tensor [^1] The trivial Killing tensor is the metric tensor $g_{\mu\nu}$ where its existence indicates the conservation of the rest mass of a moving particle in Hamiltonian systems., there are closed trajectories [Burns and Matveev, 2021].
Furthermore, as Eisenhart [Eisenhart, 1934] and Kalnins-Miller [Kalnins and Miller, 1980], [Kalnins and Miller, 1981], [Kalnins and Miller, 1983] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of valence two [Benenti, 2016]. Indeed, there is a bizarre relation between the structure of separated metrics with the structure of its characteristic Killing tensor. Benenti and Francaviglia [Benenti and Francaviglia, 1979] presented a certain example where the additional information about the metric tensor serve as a catalyst in order to obtain the structure of the Killing tensor.
Besides, there are two ways to benefit from Killing tensor, either, by assuming its existence, to find a metric or by revealing the hidden symmetries of a known metric or both. Consequently, the existence of Killing tensor in a physical problem helps us to determine a solvable system of equations through its integrability conditions (IC) and, in some cases, leads to the separation of Hamilton-Jacobi equation providing us with integrable trajectories. Regarding this, the assumption of existence of Killing tensor could serve as a promising starting point in the pursuit of “realistic” spacetimes endowed by integrable trajectories.
The main idea that motivated us to deal with the Canonical Forms of the Killing tensor was that we might find new interesting spacetimes in vacuum if we deal with more general forms of a Killing tensor with more than two distinct eigenvalues as postulated by Hauser and Malhiot in [Hauser and Malhiot, 1976]. During the last decades the only works in the literature that utilize a Killing tensor to explore new spacetimes including Hauser and Malhiot’s work on electro-vacuum [Hauser and Malhiot, 1978] and the work of one of us on interior solutions with perfect fluid [Papakostas, 1998]. Hauser-Malhiot managed to found one of the most general family of stationary axially symmetric electro-vacuum spacetimes that found independently by Carter [Carter, 1968]. But both of these works serve as the only paradigms for our research, since these are the only works both originate from the assumption of existence of a Killing tensor with two double eigenvalues.
It is worth mentioning that the Killing tensor form with two double eigenvalues is a special case of the canonical forms of a Killing tensor, and its study has yielded general and new families of exact solutions in electro-vacuum.
$K_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&0&0\\ \lambda_{1}&0&0&0\\ 0&0&0&\lambda_{2}\\ 0&0&\lambda_{2}&0\end{pmatrix}$ (Paradigm)
It is both interesting and important to explore Einstein’s spacetimes that admit more general forms of Killing tensors beyond the case of two double eigenvalues as initial premise. These more general forms represent the four canonical forms of Killing tensor. It is also intriguing to investigate whether these more general forms can lead to new solutions or generalizations of the ones already known. Additionally, we aim to investigate whether more general Petrov type solutions (e.g. Petrov type I) admit the canonical forms of the Killing tensor, as opposed to the Killing form of the Paradigm. We choose thus to take advantage of the canonical forms of Killing Tensor to discover new metrics with explicit and hidden symmetries or to acquire more information of already known metrics.
It’s also worth noting that the use of the standard metric formalism is not practical for our purpose, as it cannot accommodate the symmetries associated with a Killing tensor. In this context, the most suitable formalism is the complex vectorial formalism of spin coefficients [Cahen et al., 1967] or the Newman-Penrose formalism [Newman and Penrose, 1962]. The first reason is that during the resolution process, the classification of the gravitational field of a spacetime according to Petrov occurs in the early stages of the process. This allows us to deduce theorems of symmetries for each Petrov type [Petrov, 2000], determining an appropriate coordinate system using the Frobenius theorem of integrability.
Secondly, capitalizing on the Killing equations we obtain simplifications between the spin coefficients and additional relations to Newman-Penrose field equations. The latter plays a pivotal role in solving the field equations without the need for specific coordinate system. This is possible since the Newman-Penrose field equations are first-order differential equations of spin coefficients. In fact, the complex vectorial formalism we employ provides insights into the essential characteristics of null congruences (Shear, Divergence, Geodesic), which are related to singularity theorems.
We have attempted to establish a coherent structure for this work. In Chapter 2 we will exhibit the main points of the appointed formalism which will be revisited throughout the paper. Next, in Chapter 3 we imply a rotation around the null tetrad frame with $l^{\mu}$ fixed and we obtain the general key relations that can be applied to any canonical form except $K^{0}$. In Chapter 4 the definition of Killing tensor is given and we revisit the canonical forms of Killing tensor. After that in Chapters 5, 6 and 7 the resolution process takes place for $K^{0}$, $K^{1}$ and $K^{2,3}$, therein we present the Petrov types of the obtained solutions. The Chapter 8 before Discussion and Conclusions contains a reduced Killing form with $\lambda_{0}=0$ which is a subcase of $K^{1},K^{2},K^{3}$ Killing forms.
2 Notation of the Newman-Penrose Formalism
The Newman-Penrose Formalism is a widely known formalism that was presented by Newman and Penrose [Newman and Penrose, 1962] and was analyzed geometrically by Cahen, Debever and Defrise [Cahen et al., 1967], [Debever, 1964]. Initially, the formalism was found in order to describe the gravitational radiation in General Relativity but it was proved to have much more usefulness.
The main concept of the formalism could be briefly described as follows. The need to interpret the gravitational radiation more conveniently forces us to associate the Riemann tensor with isotropic null tetrads (light-like vectors). The latter could happen in a 3-dimensional complex bivector space ($C_{3}$) spanned by self-dual 2-forms.
2.1 Coordinate system
The metric can be put in the form
$ds^{2}=2(\boldsymbol{\theta}^{1}\boldsymbol{\theta}^{2}-\boldsymbol{\theta}^{3}\boldsymbol{\theta}^{4})$ (1)
where the general metric $g_{\mu\nu}$ is the following and equal to its inverse $g^{\mu\nu}$.
$g_{\mu\nu}=l_{\mu}n_{\nu}+n_{\mu}l_{\nu}-m_{\mu}\bar{m}_{\nu}-\bar{m}_{\mu}m_{\nu}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&-1\\ 0&0&-1&0\end{pmatrix}$ (2)
The pseudo-orthonormal basis contains two real and two complex conjugate vectors
$\boldsymbol{\theta}^{1}\equiv n_{\mu}dx^{\mu}\hskip 22.76228pt\boldsymbol{\theta}^{2}\equiv l_{\mu}dx^{\mu}\hskip 22.76228pt\boldsymbol{\theta}^{3}\equiv-\bar{m}_{\mu}dx^{\mu}\hskip 22.76228pt\boldsymbol{\theta}^{4}\equiv-m_{\mu}dx^{\mu}$ (3)
and the orthogonality properties of the vector components are the following since the rest combinations give zero.
$l_{\mu}n^{\mu}=1=-m_{\mu}\bar{m}^{\mu}$ (4)
The directional derivatives (dual basis) of the formalism are given by
$\boldsymbol{D}=l^{\mu}\partial_{\mu}\hskip 22.76228pt\boldsymbol{\Delta}=n^{\mu}\partial_{\mu}\hskip 22.76228pt\boldsymbol{\delta}=m^{\mu}\partial_{\mu}\hskip 22.76228pt\boldsymbol{\bar{\delta}}=\bar{m}^{\mu}\partial_{\mu}$
Using the Cartan’s method we can calculate the connection 1-forms ${\boldsymbol{\Gamma}^{\alpha}}_{\nu}\equiv{\Gamma^{\alpha}}_{\mu\nu}\boldsymbol{\theta}^{\mu}$.
$d\boldsymbol{\theta}^{\alpha}=-{\boldsymbol{\Gamma}^{\alpha}}_{\nu}\wedge\boldsymbol{\theta}^{\nu}$ (5)
which is explicitly written as follows
$d\theta^{1}=(\gamma+\bar{\gamma})\theta^{1}\wedge\theta^{2}+(\bar{\alpha}+\beta-\bar{\pi})\theta^{1}\wedge\theta^{3}+(\alpha+\bar{\beta}-\pi)\theta^{1}\wedge\theta^{4}-\bar{\nu}\theta^{2}\wedge\theta^{3}-\nu\theta^{2}\wedge\theta^{4}-(\mu-\bar{\mu})\theta^{3}\wedge\theta^{4}$ (6)
$d\theta^{2}=(\epsilon+\bar{\epsilon})\theta^{1}\wedge\theta^{2}+\kappa\theta^{1}\wedge\theta^{3}+\bar{\kappa}\theta^{1}\wedge\theta^{4}-(\bar{\alpha}+\beta-\tau)\theta^{2}\wedge\theta^{3}-(\alpha+\bar{\beta}-\bar{\tau})\theta^{2}\wedge\theta^{4}-(\rho-\bar{\rho})\theta^{3}\wedge\theta^{4}$ (7)
$d\theta^{3}=-(\bar{\tau}+\pi)\theta^{1}\wedge\theta^{2}-(\bar{\rho}+\epsilon-\bar{\epsilon})\theta^{1}\wedge\theta^{3}-\bar{\sigma}\theta^{1}\wedge\theta^{4}+(\mu-\gamma+\bar{\gamma})\theta^{2}\wedge\theta^{3}+\lambda\theta^{2}\wedge\theta^{4}+(\alpha-\bar{\beta})\theta^{3}\wedge\theta^{4}$ (8)
$d\theta^{4}=-(\tau+\bar{\pi})\theta^{1}\wedge\theta^{2}-\sigma\theta^{1}\wedge\theta^{3}-(\rho-\epsilon+\bar{\epsilon})\theta^{1}\wedge\theta^{4}+\bar{\lambda}\theta^{2}\wedge\theta^{3}+(\bar{\mu}+\gamma-\bar{\gamma})\theta^{2}\wedge\theta^{4}-(\bar{\alpha}-\beta)\theta^{3}\wedge\theta^{4}$ (9)
the greek letters represent the 12 complex spin coefficients. In Newman-Penrose formalism the Christoffel symbols are represented by the spin coefficients or spin connections. The relations (6)-(9) are obtained by the usage of the covariant derivatives of the null tetrads.
$n_{\mu;\alpha}=-(\epsilon+\bar{\epsilon})n_{\alpha}n_{\mu}-(\gamma+\bar{\gamma})l_{\alpha}n_{\mu}+(\alpha+\bar{\beta})m_{\alpha}n_{\mu}+(\bar{\alpha}+\beta)\bar{m}_{\alpha}n_{\mu}+\pi n_{\alpha}m_{\mu}\\ +\nu l_{\alpha}m_{\mu}-\lambda m_{\alpha}m\mu-\mu\bar{m}_{\alpha}m_{\mu}+\bar{\pi}n_{\alpha}\bar{m}_{\mu}+\bar{\nu}l_{\alpha}\bar{m}_{\mu}-\bar{\mu}m_{\alpha}\bar{m}_{\mu}-\bar{\lambda}\bar{m}_{\alpha}\bar{m}_{\mu}$ (10)
$l_{\mu;\alpha}=(\epsilon+\bar{\epsilon})n_{\alpha}l_{\mu}+(\gamma+\bar{\gamma})l_{\alpha}l_{\mu}-(\alpha+\bar{\beta})m_{\alpha}l_{\mu}-(\bar{\alpha}+\beta)\bar{m}_{\alpha}l_{\mu}-\bar{\kappa}n_{\alpha}m_{\mu}\\ -\bar{\tau}l_{\alpha}m_{\mu}+\bar{\sigma}m_{\alpha}m\mu+\bar{\rho}\bar{m}_{\alpha}m_{\mu}-\kappa n_{\alpha}\bar{m}_{\mu}-\tau l_{\alpha}\bar{m}_{\mu}+\rho m_{\alpha}\bar{m}_{\mu}+\sigma\bar{m}_{\alpha}\bar{m}_{\mu}$ (11)
$m_{\mu;\alpha}=-\kappa n_{\alpha}n_{\mu}-\tau l_{\alpha}n_{\mu}+\rho m_{\alpha}n_{\mu}+\sigma\bar{m}_{\alpha}n_{\mu}+\bar{\pi}n_{\alpha}l_{\mu}+\bar{\nu}l_{\alpha}l_{\mu}-\bar{\mu}m_{\alpha}l_{\mu}\\ -\bar{\lambda}\bar{m}_{\alpha}l_{\mu}+(\epsilon-\bar{\epsilon})n_{\alpha}m_{\mu}+(\gamma-\bar{\gamma})l_{\alpha}m_{\mu}-(\alpha-\bar{\beta})m_{\alpha}m_{\mu}+(\bar{\alpha}-\beta)\bar{m}_{\alpha}m_{\mu}$ (12)
2.2 Bivector Space
The antisymmetricity of the electromagnetic tensor $F_{\mu\nu}$ provides us with six independent components $\mu\nu=12,13,14,23,24,34$ which are considered as a bivector basis in the 6-dimensional linear space ($M_{6}$) and also as elements of the orthochronous [^2] The orthochronous Lorentz Group does not have mirroring in the timelike direction. Lorentz group $SO^{+}(1,3)$ since they generate a Lie algebra. The 6-dimensional Lie algebra SO(1,3) is isomorphic to a complex 3-dimensional Lie algebra SL(2,C). Hence, the irreducible representation of the complex Lie algebra of the Lorentz group is embodied by the self-dual bivector basis.
This basis is defined by the following relations
$\boldsymbol{Z}^{1}=\boldsymbol{\theta}^{1}\wedge\boldsymbol{\theta}^{3}=Z^{1}_{\alpha\beta}dx^{\alpha}\otimes dx^{\beta}\hskip 2.84544pt;\hskip 5.69046ptZ^{1}_{\alpha\beta}=-n_{\alpha}\bar{m}_{\beta}+n_{\beta}\bar{m}_{\alpha}$ (13)
$\boldsymbol{Z}^{2}=\boldsymbol{\theta}^{1}\wedge\boldsymbol{\theta}^{2}-\boldsymbol{\theta}^{3}\wedge\boldsymbol{\theta}^{4}=Z^{2}_{\alpha\beta}dx^{\alpha}\otimes dx^{\beta}\hskip 2.84544pt;\hskip 5.69046ptZ^{2}_{\alpha\beta}=n_{\alpha}l_{\beta}-n_{\beta}l_{\alpha}-\bar{m}_{\alpha}m_{\beta}+m_{\alpha}\bar{m}_{\beta}$ (14)
$\boldsymbol{Z}^{3}=\boldsymbol{\theta}^{4}\wedge\boldsymbol{\theta}^{2}=Z^{3}_{\alpha\beta}dx^{\alpha}\otimes dx^{\beta}\hskip 2.84544pt;\hskip 5.69046ptZ^{3}_{\alpha\beta}=-m_{\alpha}l_{\beta}+m_{\beta}l_{\alpha}$ (15)
The composite of the metric  [^3] The latin letters a,b,..,k, take values 1,2,3. in this base is
$\gamma^{ab}=4\left[{\delta^{a}}_{(1}{\delta^{b}}_{3)}-{\delta^{a}}_{2}{\delta^{b}}_{2}\right]=\begin{pmatrix}0&0&2\\ 0&-4&0\\ 2&0&0\end{pmatrix}$ (16)
The complex connection 1-forms $\boldsymbol{\sigma}_{b}^{a}$ is produced by derivation of the basis $\boldsymbol{Z}^{a}$, i.e
$d\boldsymbol{Z}^{a}=-\boldsymbol{\sigma}^{a}_{b}\wedge\boldsymbol{Z}^{b}$ (17)
the vectorial connection 1-form $\boldsymbol{\sigma}_{a}$ is defined by
$\boldsymbol{\sigma}^{a}_{b}=8\epsilon^{kac}\boldsymbol{\sigma}_{k}\gamma_{cb}\hskip 5.69046pt\Leftrightarrow\hskip 5.69046pt\boldsymbol{\sigma}_{k}=\frac{1}{8}\epsilon_{kac}\gamma^{cb}\boldsymbol{\sigma}^{a}_{b}=\kappa_{k\mu}\boldsymbol{\theta}^{\mu}$ (18)
where $\epsilon_{abc}$ is the Levi-Civita tensor and the tetrad components $\kappa_{k\mu}$ contain the 12 complex spin coefficients
$\kappa_{k\mu}=\begin{bmatrix}\kappa&\tau&\sigma&\rho\\ \epsilon&\gamma&\beta&\alpha\\ \pi&\nu&\mu&\lambda\\ \end{bmatrix}$ (19)
The complex curvature 2-forms $\boldsymbol{\Sigma}^{b}_{d}$ are defined by
[…]
The corresponding expanding of $\boldsymbol{\Sigma}_{a}$ with respect to the basis of $\left(\boldsymbol{Z}^{a},\boldsymbol{\bar{Z}}^{a}\right)$ is given by
$\boldsymbol{\Sigma}_{a}=(C_{ab}-\frac{1}{6}R\gamma_{ab})\boldsymbol{Z}^{b}+E_{a\bar{b}}\boldsymbol{\bar{Z}}^{b},$ (21)
where these quantities are related with the curvature components of the formalism
$C_{ab}=\begin{bmatrix}\Psi_{0}&\Psi_{1}&\Psi_{2}\\ \Psi_{1}&\Psi_{2}&\Psi_{3}\\ \Psi_{2}&\Psi_{3}&\Psi_{4}\\ \end{bmatrix},\hskip 28.45274ptE_{a\bar{b}}=\begin{bmatrix}\Phi_{00}&\Phi_{01}&\Phi_{02}\\ \Phi_{10}&\Phi_{11}&\Phi_{12}\\ \Phi_{20}&\Phi_{21}&\Phi_{22}\\ \end{bmatrix}$ (22)
In this formalism, the 10 Weyl’s components are represented by the 5 complex scalar functions.
$\Psi_{0}=C_{\kappa\lambda\mu\nu}l^{\kappa}m^{\lambda}l^{\mu}m^{\nu}$
$\Psi_{1}=C_{\kappa\lambda\mu\nu}l^{\kappa}n^{\lambda}l^{\mu}m^{\nu}$
$\Psi_{2}=\frac{1}{2}C_{\kappa\lambda\mu\nu}l^{\kappa}n^{\lambda}\left[l^{\mu}n^{\nu}-m^{\mu}\bar{m}^{\nu}\right]$ (23)
$\Psi_{3}=C_{\kappa\lambda\mu\nu}n^{\kappa}l^{\lambda}n^{\mu}\bar{m}^{\nu}$
$\Psi_{4}=C_{\kappa\lambda\mu\nu}n^{\kappa}\bar{m}^{\lambda}n^{\mu}\bar{m}^{\nu}$
The Ricci tensor components are represented by $E_{a\bar{b}}$ and they are divided in to real and complex components. All these quantities describe the main parts of the EFE. The EFE in this formalism are represented by the corresponding field equations which are known either as Newman-Penrose Equations (NPE) or as Ricci identities [Newman and Penrose, 1962].
$D\rho-\bar{\delta}\kappa={\rho}^{2}+\sigma\bar{\sigma}+\rho(\epsilon+\bar{\epsilon})-\bar{\kappa}\tau-\kappa\left[2(\alpha+\bar{\beta})+(\alpha-\bar{\beta})-\pi\right]$ (a)
$\delta\kappa-D\sigma=-(\rho+\bar{\rho}+3\epsilon-\bar{\epsilon})\sigma+\kappa\left[\tau-\bar{\pi}+2(\bar{\alpha}+\beta)-(\bar{\alpha}-\beta)\right]-\Psi_{o}$ (b)
$D\tau=\Delta\kappa+\rho(\tau+\bar{\pi})+\sigma(\pi+\bar{\tau})+\tau(\epsilon-\bar{\epsilon})-2\kappa\gamma-\kappa(\gamma+\bar{\gamma})+\Psi_{1}$ (c)
$D\nu-\Delta\pi=\mu(\pi+\bar{\tau})+\lambda(\bar{\pi}+\tau)+\pi(\gamma-\bar{\gamma})-2\nu\epsilon-\nu(\epsilon+\bar{\epsilon})+\Psi_{3}$ (i)
$\bar{\delta}\pi-D\lambda=-\pi(\pi+\alpha-\bar{\beta})-\bar{\sigma}\mu+\nu\bar{\kappa}+\lambda(3\epsilon-\bar{\epsilon})$ (g)
$\delta\tau-\Delta\sigma=\mu\sigma+\bar{\lambda}\rho+\tau(\tau-\bar{\alpha}+\beta)-\sigma(3\gamma-\bar{\gamma})-\bar{\nu}\kappa$ (p)
$D\mu-\delta\pi=\mu\bar{\rho}+\sigma\lambda+\pi(\bar{\pi}-\bar{\alpha}+\beta)-\mu(\epsilon+\bar{\epsilon})-\kappa\nu+\Psi_{2}+2\Lambda$ (h)
$\delta\nu-\Delta\mu=\mu(\mu+\gamma+\bar{\gamma})+\lambda\bar{\lambda}-\bar{\nu}\pi+\nu(\tau-2(\bar{\alpha}+\beta)+(\bar{\alpha}-\beta))$ (n)
$\Delta\rho-\bar{\delta}\tau=-(\bar{\mu}\rho+\sigma\lambda)-\tau(\bar{\tau}+\alpha-\bar{\beta})+\nu\kappa+\rho(\gamma+\bar{\gamma})-\Psi_{2}-2\Lambda$ (q)
$\delta\rho-\bar{\delta}\sigma=\rho(\bar{\alpha}+\beta)-\sigma(3\alpha-\bar{\beta})+\tau(\rho-\bar{\rho})+\kappa(\mu-\bar{\mu})-\Psi_{1}$ (k)
$\bar{\delta}\mu-\delta\lambda=-\mu(\alpha+\bar{\beta})-\pi(\mu-\bar{\mu})-\nu(\rho-\bar{\rho})-\lambda(\bar{\alpha}-3\beta)+\Psi_{3}$ (m)
$D\alpha-\bar{\delta}\epsilon=\alpha(\rho+\bar{\epsilon}-2\epsilon)+\beta\bar{\sigma}-\bar{\beta}\epsilon-\kappa\lambda-\bar{\kappa}\gamma+\pi(\epsilon+\rho)$ (d)
$D\beta-\delta{\epsilon}=\sigma(\alpha+\pi)+\beta(\bar{\rho}-\bar{\epsilon})-\kappa(\mu+\gamma)-\epsilon(\bar{\alpha}-\bar{\pi})+\Psi_{1}$ (e)
$\Delta\alpha-\bar{\delta}\gamma=\nu(\epsilon+\rho)-\lambda(\tau+\beta)+\alpha(\bar{\gamma}-\bar{\mu})+\gamma(\bar{\beta}-\bar{\tau})-\Psi_{3}$ (r)
$-\Delta\beta+\delta\gamma=\gamma(\tau-\bar{\alpha}-\beta)+\mu\tau-\sigma\nu-\epsilon\bar{\nu}-\beta(\gamma-\bar{\gamma}-\mu)$ (o)
$\delta\alpha-\bar{\delta}\beta=\mu\rho-\sigma\lambda+\alpha(\bar{\alpha}-\beta)-\beta(\alpha-\bar{\beta})+\gamma(\rho-\bar{\rho})+\epsilon(\mu-\bar{\mu})-\Psi_{2}+\Lambda$ (l)
$D\gamma-\Delta\epsilon=\alpha(\tau+\bar{\pi})+\beta(\bar{\tau}+\pi)-\gamma(\epsilon+\bar{\epsilon})-\epsilon(\gamma+\bar{\gamma})+\Psi_{2}-\Lambda-\kappa\nu+\tau\pi$ (f)
$\bar{\delta}\nu-\Delta\lambda=\lambda(\mu+\bar{\mu}+3\gamma-\bar{\gamma})-\nu\left[2(\alpha+\bar{\beta})+(\alpha-\bar{\beta})+\pi-\bar{\tau}\right]+\Psi_{4}$ (j)
The Bianchi Identities (BI) are given by the following relations
$\bar{\delta}\Psi_{0}-D\Psi_{1}=(4\alpha-\pi)\Psi_{0}-2(2\rho+\epsilon)\Psi_{1}+3\kappa\Psi_{2}$ (I)
$\bar{\delta}\Psi_{1}-D\Psi_{2}=\lambda\Psi_{0}+2(\alpha-\pi)\Psi_{1}-3\rho\Psi_{2}+2\kappa\Psi_{3}$ (II)
$\bar{\delta}\Psi_{2}-D\Psi_{3}=-3\pi\Psi_{2}+2\lambda\Psi_{1}+2(\epsilon-\rho)\Psi_{3}+\kappa\Psi_{4}$ (III)
$\bar{\delta}\Psi_{3}-D\Psi_{4}=-2(\alpha+2\pi)\Psi_{3}+(4\epsilon-\rho)\Psi_{4}+3\lambda\Psi_{2}$ (IV)
$\Delta\Psi_{0}-\delta\Psi_{1}=(4\gamma-\mu)\Psi_{0}-2(2\tau+\beta)\Psi_{1}+3\sigma\Psi_{2}$ (V)
$\Delta\Psi_{1}-\delta\Psi_{2}=\nu\Psi_{0}+2(\gamma-\mu)\Psi_{1}-3\tau\Psi_{2}+2\sigma\Psi_{3}$ (VI)
$\Delta\Psi_{2}-\delta\Psi_{3}=\sigma\Psi_{4}+2\nu\Psi_{1}-3\mu\Psi_{2}+2(\beta-\tau)\Psi_{3}$ (VII)
$\Delta\Psi_{3}-\delta\Psi_{4}=3\nu\Psi_{2}-2(\gamma+2\mu)\Psi_{3}+(4\beta-\tau)\Psi_{4}$ (VIII)
Also, the Lie bracket plays an important role to the theory, since the commutation relations emerged by its implication on the vectors $n^{\mu},l^{\mu},m^{\mu},\bar{m}^{\mu}$. The proper definition reads as follows for an arbitrary vector basis.
$[\boldsymbol{e}_{\mu},\boldsymbol{e}_{\nu}]=-2{\Gamma^{\sigma}}_{[\mu\nu]}\boldsymbol{e}_{\sigma}$ (24)
The commutations relations (CR) of the theory are given by
$[n^{\mu},l^{\mu}]=[D,\Delta]=(\gamma+\bar{\gamma})D+(\epsilon+\bar{\epsilon})\Delta-(\pi+\bar{\tau})\delta-(\bar{\pi}+\tau)\bar{\delta}$ (CR1)
$[(\delta+\bar{\delta}),D]=(\alpha+\bar{\alpha}+\beta+\bar{\beta}-\pi-\bar{\pi})D+(\kappa+\bar{\kappa})\Delta-(\bar{\sigma}+\bar{\rho}+\epsilon-\bar{\epsilon})\delta-(\sigma+\rho-\epsilon+\bar{\epsilon})\bar{\delta}$ ($CR2_{+}$)
$[(\delta-\bar{\delta}),D]=(-\alpha+\bar{\alpha}+\beta-\bar{\beta}+\pi-\bar{\pi})D+(\kappa-\bar{\kappa})\Delta-(\bar{\rho}-\bar{\sigma}+\epsilon-\bar{\epsilon})\delta+(\rho-\sigma-\epsilon+\bar{\epsilon})\bar{\delta}$ ($CR2_{-}$)
$[(\delta+\bar{\delta}),\Delta]=-(\nu+\bar{\nu})D+(\tau+\bar{\tau}-\alpha-\bar{\alpha}-\beta-\bar{\beta})\Delta+(\mu+\lambda-\gamma+\bar{\gamma})\delta+(\bar{\mu}+\bar{\lambda}+\gamma-\bar{\gamma})\bar{\delta}$ ($CR3_{+}$)
$[(\delta-\bar{\delta}),\Delta]=-(\nu-\bar{\nu})D+(\tau-\bar{\tau}+\alpha-\bar{\alpha}-\beta+\bar{\beta})\Delta+(\mu-\lambda-\gamma+\bar{\gamma})\delta-(\bar{\mu}-\bar{\lambda}+\gamma-\bar{\gamma})\bar{\delta}$ ($CR3_{-}$)
$[\delta,\bar{\delta}]=-(\mu-\bar{\mu})D-(\rho-\bar{\rho})\Delta+(\alpha-\bar{\beta})\delta-(\bar{\alpha}-\beta)\bar{\delta}$ ($CR4$)
All the above sets of equations contribute to the NPE, the BI and the CR of the basis vectors. Despite the fact that we have to solve a considerably larger number of equations this formalism has great advantages. Gauge transformations of the tetrad can be used to simplify the field equations and we can easily extract invariant properties of the gravitational field (Petrov types) without using a coordinate basis [Stephani et al., 2009]. Also, it allows us to search for solutions with specific special features, such as the presence of one or two null directions that might be singled out by physical or geometric considerations.
3 Null rotation
The IC along with the NPE end up to be a cumbersome system of equations, thus, there are various approaches to obtain simplifications for our problem. The most used method involves implementing transformations by applying a rotation within the null tetrad frame (Null rotation) or exploring different options among the spin coefficients determining the rotation parameters (Lorentz transformations).
In this work we choose to take advantage of the conformal symmetry of a rotation around one of the real null vectors, namely we choose $l^{\mu}$ to be fixed, although we will exhibit both transformations in the relations below, wherein the exponentials correspond to Lorentz transformations and the parentheses correspond to null rotation [Stewart, 1993]. We define the complex rotation parameters $t\equiv a+ib$ and $p\equiv c+id$.
$\tilde{\theta}^{1}=e^{-a}(\theta^{1}+p\bar{p}\theta^{2}+\bar{p}\theta^{3}+p\theta^{4})=\tilde{n}_{\mu}dx^{\mu}$
$\tilde{\theta}^{2}=e^{a}\theta^{2}=\tilde{l}_{\mu}dx^{\mu}$
$\tilde{\theta}^{3}=e^{-ib}(\theta^{3}+p\theta^{2})=-\tilde{\bar{m}}_{\mu}dx^{\mu}$
$\tilde{\theta}^{4}=e^{ib}(\theta^{4}+\bar{p}\theta^{2})=-\tilde{m}_{\mu}dx^{\mu}$
Subsequently, we require that the Killing tensor remains invariant under rotation. For instance we present the following compact form where $q=0,\pm 1$ for $K^{1}$ and $K^{2},K^{3}$ accordingly because the null rotation is not applicable for $K^{0}$ form.
$K^{1,2,3}=\lambda_{0}(\tilde{\theta}^{1}\otimes\tilde{\theta}^{1}+q\tilde{\theta}^{2}\otimes\tilde{\theta}^{2})+\lambda_{1}(\tilde{\theta}^{1}\otimes\tilde{\theta}^{2}+\tilde{\theta}^{2}\otimes\tilde{\theta}^{1})+\lambda_{2}(\tilde{\theta}^{3}\otimes\tilde{\theta}^{4}+\tilde{\theta}^{4}\otimes\tilde{\theta}^{3})+\lambda_{7}(\tilde{\theta}^{3}\otimes\tilde{\theta}^{3}+\tilde{\theta}^{4}\otimes\tilde{\theta}^{4})$
It is easy for someone to prove that the only non-zero rotation parameter is $t$ in case where the diagonal elements of the tensor are absent. This is valid due to the existence of the cross terms $\tilde{\theta}^{1}\otimes\tilde{\theta}^{2}$ and $\tilde{\theta}^{3}\otimes\tilde{\theta}^{4}$. Hence, in the next chapters scoping to obtain simplifications for our spin coefficients we may annihilate either $\lambda_{0}$ or $\lambda_{7}$ since the absence of other elements like $\lambda_{1}$ or $\lambda_{2}$ do not contribute to our cause. Let’s proceed with the implication of the rotation on the spin coefficients. As we observe, the only further transformations that can be applied are Lorentz transformations since $p$ is equal to zero due to the invariance of the Killing tensor. Along these lines, to achieve the most general simplification, we won’t correlate the spin coefficients with each other while determining the non-zero rotation parameter $t$. Thus, the transformation we apply is solely the null rotation.
$\tilde{\tau}=e^{ib}\tau\hskip 28.45274pt\tilde{\pi}=e^{-ib}\pi$
$\tilde{\rho}=e^{a}\rho\hskip 28.45274pt\tilde{\mu}=e^{-a}\mu$
$\tilde{\kappa}=e^{2a+ib}\kappa\hskip 28.45274pt\tilde{\nu}=e^{-(2a+ib)}\nu$
$\tilde{\sigma}=e^{a+2ib}\sigma\hskip 28.45274pt\tilde{\lambda}=e^{-(a+2ib)}\lambda$
$\tilde{\beta}=e^{ib}\left(\beta+\frac{\delta t}{2}\right)\hskip 28.45274pt\tilde{\alpha}=e^{-ib}\left(\alpha+\frac{\bar{\delta}{t}}{2}\right)$
$\tilde{\epsilon}=e^{a}\left(\epsilon+\frac{Dt}{2}\right)\hskip 28.45274pt\tilde{\gamma}=e^{-a}\left(\gamma+\frac{\Delta t}{2}\right)$
There are two different kinds of simplifications that can be acquired by the capitalization of the annihilation of the tilded spin coefficients $\tilde{\epsilon},\tilde{\gamma},\tilde{\alpha},\tilde{\beta}$. The simplest simplification emerges by the correlation of the spin coefficient with the derivative of the rotation parameter $t$. In case where $\lambda_{0}=0$ the rotation parameter becomes $t=a$, differently, in case where $\lambda_{7}$ the non-zero rotation parameter is $t=ib$. The latter has significant impact on the spin coefficients. When $t=a$ the spin coefficients $\epsilon,\gamma,\alpha,\beta$ depends on the directional derivatives of real quantities yielding the following relations
$\epsilon-\bar{\epsilon}=0$
$\gamma-\bar{\gamma}=0$
$\alpha-\bar{\beta}=0$
On the other hand when $t=ib$ we get
$\epsilon+\bar{\epsilon}=0$
$\gamma+\bar{\gamma}=0$
$\alpha+\bar{\beta}=0$
The second kind of simplification takes place when we substitute the four spin relations into the CR and we compare the outcome with the NPE (f),(l),(e),(o) resulting in the key relations. The latter unfolds the branches of the solutions. We postulate the most general case of the obtained relations after the comparison with NPE.
$\Psi_{2}-\Lambda=\kappa\nu-\tau\pi$ (i)
$\Psi_{1}=\kappa\mu-\sigma\pi$ (ii)
$\Psi_{2}-\Lambda=\mu\rho-\sigma\lambda$ (iii)
$\mu\tau-\sigma\nu=0$ (iv)
The null rotation provides us with these useful relations which connect the spin coefficients along with the Weyl components. This result depends mainly on the form of the Killing Tensor, since we demand the preservation of Killing tensor. Essentially, the lack either of $\lambda_{0}$ or $\lambda_{7}$ allows us to obtain simplifications such as the key relations  [^4] The annihilation of $\lambda_{0}$ serves as a possible reduction to this canonical form although the absence of $\lambda_{0}$ in $K^{1}$, $K^{2}$, $K^{3}$ makes them to coincide. This choice is exhibited in Chapter 9..
Remark The application of null rotation and Lorentz transformations are transformative processes that yield valuable relationships connecting not only the spin coefficients amongst themselves but also with the Weyl components through the commutation relations, once the tilded spin coefficients have been annihilated. The outcome critically hinges on the preservation of the structure of the Killing tensor and the form of the spin coefficients. Importantly, the absence of either $\lambda_{0}$ or $\lambda_{7}$ is the catalyst for these relationships. Employing the complete form of the Killing tensor $K^{1,2,3}_{\mu\nu}$, conversely, provides no insight into this implied symmetry since both $t$ and $p$ are nullified.
4 The canonical forms of Killing tensor
The preservation of geometry of spacetime through a transformation reveals the existence of symmetries. These symmetries leave invariant the elements which characterize the geometry, that is referred to the metric tensor and, by extension, the action. Thus, the consideration of symmetries in the resolution process of the Einstein’s equations is indispensable. The non-linear character of the equations of gravity obligates us to introduce further information through symmetries to obtain a solvable-overdetermined system of equations.
4.1 Killing Tensor
The geodesic flow is a Hamiltonian system on the cotangent bundle
$H=\frac{1}{2}g^{\mu\nu}p_{\mu}p_{\nu}$ (25)
where $p_{\mu}$ are the coordinates on the cotangent spaces or equivalently the canonical momenta of an observer. Then, an integral of motion could be defined as follows [Kruglikov and Matveev, 2016], [Sommers, 1973].
[…]
The function $\boldsymbol{\mathcal{K}}$: $T^{*}M\rightarrow\mathbb{R}$ is called polynomial of momenta and is defined as
$\boldsymbol{\mathcal{K}}(x,p)\equiv K^{\mu}p_{\mu}+K^{\mu\nu}p_{\mu}p_{\nu}+K^{\mu\nu\sigma}p_{\mu}p_{\nu}p_{\sigma}+K^{\mu\nu\sigma\rho}p_{\mu}p_{\nu}p_{\sigma}p_{\rho}$ (27)
[…]
$K_{(\mu;\nu)}=0$
$K_{(\mu\nu;\alpha)}=0$
$K_{(\mu\nu\sigma;\alpha)}=0$
$K_{(\mu\nu\sigma\rho;\alpha)}=0$
Killing tensors of rank r give rise to a homogeneous constant of motion of degree r in momenta. The inhomogeneous polynomial integrals of geodesic motion can be decomposed to their homogeneous parts and also to the corresponding parts that are associated with the Killing tensors with the equivalent rank. The Killing tensor of rank 1 equals to Killing vector and generates continuous symmetry transformations, these symmetries are called explicit [Garfinkle and Glass, 2010]. The symmetries that correspond to higher-order ranks of momenta associated with Killing tensors of rank ($r>1$) are called hidden symmetries [Frolov et al., 2017]. The investigation of special objects such Killing tensor or Killing-Yano is basically a devilish way to peep into the phase space searching for hidden symmetries [Krtouš et al., 2007]. Indeed, the explicit symmetries in a Hamiltonian system always could be “dragged up” instead of hidden symmetries. These kinds of symmetries emerge during the study of the dynamics of a system featuring the conserved quantities of the system or one-parameter isometries which is equivalent with the admission of existence of Killing vectors.
As Eisenhart [Eisenhart, 1934] and Kalnins-Miller [Kalnins and Miller, 1980], [Kalnins and Miller, 1981], [Kalnins and Miller, 1983] showed, the geodesic separation is correlated with the existence of Killing vectors and Killing tensors of order two [Benenti, 2016]. Indeed, there is a bizarre relation between the structure of separated metrics with the structure of its characteristic Killing tensor. Benenti and Francaviglia [Benenti and Francaviglia, 1979] present a certain example where the additional information about the metric tensor serve as a catalyst in order to obtain the structure of the Killing tensor.
Let us proceed with the definition of Killing tensor: Any symmetric tensor of order 2 whose the symmetric part of his covariant derivative vanishes is called a Killing tensor.
$K_{(\mu\nu;\alpha)}=0$
The trivial Killing tensor is the metric tensor $g_{\mu\nu}$ where its existence indicates the conservation of the rest mass of a moving particle in Hamiltonian systems. The Hamiltonian is a conserved quantity of the problem since it is correlated with the conserved rest mass.
$\mathcal{H}=\frac{\bar{m}^{2}}{2}=\frac{1}{2}g_{\mu\nu}u^{\mu}u^{\nu}$ (28)
At last, the usage of canonical forms of a Killing tensor could be proved fruitful since can be used as a starter culture in order to discover spacetimes with hidden symmetries.
4.2 Canonical forms
Obtaining the canonical forms of a symmetric 2nd-rank tensor can be a challenging task when approached algebraically. However, in a symmetric matrix the additional symmetries could aid us to find its canonical forms geometrically. The canonical forms contain the minimum number of independent scalars which compose the eigenvalues during the diagonalization [Churchill, 1932]. The equation for diagonalization in any tensor takes the following form.
${K^{\mu}}_{\nu}z^{\nu}-\lambda z^{\mu}=0$ (29)
In this manner, it becomes clear that the operation of our tensor or our linear vector function on a vector leaves it unchanged. In this context, it is known that every 2nd-order symmetric tensor defines a linear mapping that transforms a vector k into another vector v unless the vector is an eigenvector. We are searching for these types of directions, which, of course, do not alter the quadratic form of the metric. In our formalism, the norm of a vector must remain invariant.
$\textbf{x}\cdot\textbf{x}=x^{\mu}g_{\mu\nu}x^{\nu}=2(x^{1}x^{2}-x^{3}x^{4})$ (30)
Algebraically, the result is that a symmetric matrix $K$ is reducible under an orthogonal transformation with a matrix P to a canonical form $PKP^{-1}$ in which all non-diagonal elements are zero.
A similar analysis can also be found in Landau’s book on p. 271 regarding the stress-energy-momentum tensor [Landau and Lifschitz, 1975]. In this section, he mentions that this procedure is, in fact, the application of the “Principal Axis Theorem” or “Spectral Theorem” for a matrix. However, it appears that the author applies diagonalization to a symmetric tensor in a covariant form. This approach is not correct in general. Diagonalization should be applied to a tensor in a mixed tensor form, as shown in equation (29). This is always valid because a term proportional to the metric merely shifts all eigenvalues by the same amount, the latter is also denoted in [Stephani et al., 2009].
4.2.1 The presence of null vectors within planes
In line with the work of Churchill we obtained the Canonical forms of Killing tensor [Churchill, 1932]. The study of Churchill was operated in pseudo-Euclidean spacetime with signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and the object of his study was a linear symmetric vector function whose components are represented by a symmetric tensor of valence 2. In this reference Churchill was based on the work of Rainich who studied the antisymmetricity of the electromagnetic tensor [Rainich, 1925] and he remarks that: “It is known that every linear vector function in four-dimensional space has at least one invariable plane”.
Due to the last statement Churchill operated his calculation with a specific manner. He lied his $k,l$ vectors (the corresponding $m,\bar{m}$ of our formalism) in the invariable plane which does not contain any null vectors. With this choice all the canonical forms has the same downward-right block where there are two real distinct eigenvalues. Hence, all forms have different eigenvalues only in the upward-left block where the existence of null vectors takes place.
In the following segment we present the classification of canonical forms, therein the statement of the previous paragraph is evident since all the canonical forms proved to have the same eigenvalues $-(\lambda_{2}\pm\lambda_{7})$ in the 2-dimensional plane where non-null vectors are present. Although, the $K^{0}$ form serves as an exception since it has one triple eigenvalue with multiplicity 1. Due to this necessary condition one can prove that the triple eigenvalue satisfies either the first relation
$\lambda_{1}=-(\lambda_{2}+\lambda_{7})$
or the second one.
$\lambda_{1}=-(\lambda_{2}-\lambda_{7})$
After this introduction let us proceed to the following classification. Our pseudo-Euclidean spacetime comprises three types of planes [Hall, 1976]. It was helpful to mention that our framework is described by a null tetrad frame, hence, our forms have a totally different shape. Besides, we categorize our cases differently. In our Case 0, there is a plane with one null vector and it is characterized as a singular case. In Case 1 and Case 3, there are two null vectors lie within the plane, and in Case 2, there are no null vectors within our plane [^5] For this case we expect that the two blocks upward-left and downward-right have the exact same form with different components. [Rainich, 1952].
In Case 0 there is only one null vector lies within the plane, using this we obtain the $K^{0}$ form. Using the diagonalization procedure one can easily find that it has one triple eigenvalue and it is a Jordan canonical form. The diagonalization reveals that in order to have one timelike and three spacelike eigenvectors as a necessary condition for our canonical form there are two cases for the triple eigenvalue as we mentioned few lines above. Thus, the canonical form for $K^{0}$ results to
${K{{}^{0}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&-p&-\bar{p}\\ \lambda_{1}&0&0&0\\ -p&0&\lambda_{7}&\lambda_{2}\\ -\bar{p}&0&\lambda_{2}&\lambda_{7}\end{pmatrix}\hskip 8.5359ptp=-\bar{p}=\pm 1$ (31)
In Case 1 we have two canonical forms with one double eigenvalue called as $K^{1}$. This case produces two forms but considered as one by us. The reason is based on the symmetry of the symmetrical null tetrad frame. This symmetry referred to the interchanges between the tetrads $n^{\mu}\leftrightarrow l^{\mu}$ and $m^{\mu}\leftrightarrow\bar{m}^{\mu}$. Regarding this, even if the spin coefficients interchange the result remains the same.
${K{{}^{1a}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}\hskip 2.84544pt{K{{}^{1b}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&0&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (32)
Case 2 contains also a canonical form with 4 distinct real eigenvalues called as $K^{2}$.
${K{{}^{2}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (33)
Finally, when considering the $K^{3}$ form, it possesses a pair of eigenvalues that are complex conjugates. In a more general scenario, if the form have complex eigenvalues, the tensor can be diagonalized in terms of complex pairs, where each pair consists of eigenvalues that are complex conjugates of each other. However, to maintain the requirement of having one timelike and three spacelike eigenvalues, our tensor can only admit one pair of complex conjugate eigenvalues [Landau and Lifschitz, 1975].
${K{{}^{3}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&-\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (34)
It should be noted that the only difference between $K^{2}$ and $K^{3}$ could be described via a factor $q$. We choose to deal simultaneously with forms $K^{2}$ and $K^{3}$ with the usage of the parameter $q=\pm 1$ that gives us the 2nd and 3rd forms for $+1$ and $-1$ accordingly.
4.2.2 The diagonalized form of the Canonical forms
We present the diagronalized canonical forms of the Killing tensor.
${K^{0\mu}}_{\nu}=\begin{cases}\lambda_{1}=-(\lambda_{2}+\lambda_{7})\ ;\ \begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&1&0\\ 0&0&\lambda_{1}&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}\\ \lambda_{1}=-(\lambda_{2}-\lambda_{7})\ ;\ \begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&1&0\\ 0&0&\lambda_{1}&0\\ 0&0&0&-(\lambda_{2}+\lambda_{7})\end{pmatrix}\end{cases}$ (Case 0)
${K^{1\mu}}_{\nu}=\begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 1)
${K^{2\mu}}_{\nu}=\begin{pmatrix}\lambda_{0}+\lambda_{1}&0&0\\ 0&\lambda_{0}-\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 2)
${K{{}^{3\mu}}}_{\nu}=\begin{pmatrix}\lambda_{0}+i\lambda_{1}&0&0\\ 0&\lambda_{0}-i\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 3)
In conclusion, the canonical forms encompass all the needed information of an arbitrary Killing tensor, categorized according to the number of its eigenvalues and their multiplicities. It is also obvious that the four canonical forms are generalizations of the Killing form with two double eigenvalues (Paradigm). The solution extraction through the canonical forms scoping to elevate the usage of Killing tensor.
5 Solutions of $K^{0}_{\mu\nu}$ form
We initiate the second part of this work by the study of the first canonical form $K^{0}_{\mu\nu}$ form.
$K^{0}_{\mu\nu}=\lambda_{1}(l_{\mu}n_{\nu}+n_{\mu}l_{\nu})-p(n_{\mu}m_{\nu}+m_{\mu}n_{\nu})-\bar{p}(n_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}n_{\nu})\\ +\lambda_{7}(m_{\mu}m_{\nu}+\bar{m}_{\mu}\bar{m}_{\nu})+\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu});\hskip 5.69046ptp=-\bar{p}=\pm 1$ (35)
It will be proved really helpful to define the factor $Q$, which is a real quantity since it depends on the elements of each canonical form. In the next chapters the corresponding factor will be redefined for each form differently. The factor $Q$ for $K^{0}_{\mu\nu}$ is defined by
$Q\equiv\frac{\lambda_{7}}{\lambda_{1}+\lambda_{2}}$ (36)
The relations (10)-(12) are indispensable in order to extract any useful relation by the Killing equation $K^{0}_{(\mu\nu;\alpha)}=0$. The annihilation of the symmetrised covariant derivation of our canonical form yields the following relations.
$\nu(\lambda_{1}+\lambda_{2})+\bar{\nu}\lambda_{7}=0$ (37)
$\kappa p+\kappa\bar{p}=0$ (38)
$\bar{\kappa}(\lambda_{1}+\lambda_{2})+\kappa\lambda_{7}+(\rho-\bar{\epsilon})p+\bar{\sigma}\bar{p}=0$ (39)
$\kappa(\lambda_{1}+\lambda_{2})+\bar{\kappa}\lambda_{7}+(\bar{\rho}-\epsilon)p+\sigma\bar{p}$ (40)
$D\lambda_{1}+(\tau-\bar{\pi})p+(\bar{\tau}-\pi)\bar{p}=0$ (41)
$\Delta\lambda_{1}-\bar{\nu}p-\nu\bar{p}=0$ (42)
$-\delta\lambda_{1}+(\bar{\pi}-\tau)(\lambda_{1}+\lambda_{2})+(\pi-\bar{\tau})\lambda_{7}+\bar{\lambda}p+(\mu+2\gamma)\bar{p}=0$ (43)
$-\bar{\delta}\lambda_{1}+(\pi-\bar{\tau})(\lambda_{1}+\lambda_{2})+(\bar{\pi}-\tau)\lambda_{7}+(\bar{\mu}+2\bar{\gamma})p+\lambda\bar{p}=0$ (44)
$D\lambda_{2}+(\rho+\bar{\rho})(\lambda_{1}+\lambda_{2})+(\sigma+\bar{\sigma})\lambda_{7}-(\bar{\pi}+2\bar{\alpha})p-(\pi+2\alpha)\bar{p}=0$ (45)
$\Delta\lambda_{2}-(\mu+\bar{\mu})(\lambda_{1}+\lambda_{2})-(\lambda+\bar{\lambda})(\lambda_{1}+\bar{\lambda})-\bar{\nu}p-\nu\bar{p}=0$ (46)
$-2\delta\lambda_{2}-\bar{\delta}\lambda_{7}+2(\alpha-\bar{\beta})\lambda_{7}+2\bar{\lambda}p+2(\mu+\bar{\mu})\bar{p}=0$ (47)
$-2\bar{\delta}\lambda_{2}-\delta\lambda_{7}+2(\bar{\alpha}-\beta)\lambda_{7}+2(\mu+\bar{\mu})p+2\lambda\bar{p}=0$ (48)
$D\lambda_{7}+2\bar{\sigma}(\lambda_{1}+\lambda_{2})+2(\rho+\epsilon-\bar{\epsilon})\lambda_{7}-2p(\pi+2\bar{\beta})=0$ (49)
$D\lambda_{7}+2\sigma(\lambda_{1}+\lambda_{2})+2(\bar{\rho}-\epsilon+\bar{\epsilon})\lambda_{7}-2\bar{p}(\bar{\pi}+2\beta)=0$ (50)
$\Delta\lambda_{7}-2\lambda(\lambda_{1}+\lambda_{2})+2(\gamma-\bar{\gamma}-\bar{\mu})\lambda_{7}-2\nu p=0$ (51)
$\Delta\lambda_{7}-2\bar{\lambda}(\lambda_{1}+\lambda_{2})+2(\bar{\gamma}-\gamma-\mu)\lambda_{7}-2\bar{\nu}\bar{p}=0$ (52)
$-\delta\lambda_{7}-2(\bar{\alpha}-\beta)\lambda_{7}+2\bar{p}\bar{\lambda}=0$ (53)
$-\bar{\delta}\lambda_{7}-2(\alpha-\bar{\beta})\lambda_{7}+2p\lambda=0$ (54)
We ought to mark that the annihilation of the real and the imaginary part of the equation (37) yields the following two possibilities.
$(\nu+\bar{\nu})(\lambda_{1}+\lambda_{2}+\lambda_{7})=0\Leftrightarrow(\nu+\bar{\nu})(Q+1)=0$
$(\nu-\bar{\nu})(\lambda_{1}+\lambda_{2}-\lambda_{7})=0\Leftrightarrow(\nu-\bar{\nu})(Q-1)=0$
We are already aware that $Q+1=0$ or $Q-1=0$. These two possibilities are the necessary conditions to the existence of a 3-dimensional Jordan block and also guarantee the existence of one timelike and three spacelike eigenvectors. Let us proceed with the first choice where $Q+1=0$. Then, the equations regardless the value of the factor $p$ yield
$\epsilon-\bar{\epsilon}=\kappa-\bar{\kappa}=\nu=\gamma=\mu=\lambda=0$ (55)
$\pi-\bar{\pi}+2(\alpha-\bar{\alpha})=0$ (56)
$\pi+2\bar{\beta}=0$ (57)
$\rho-\bar{\sigma}=\epsilon$ (58)
$\pi-\bar{\pi}+\tau-\bar{\tau}=0$ (59)
$D\lambda_{1}=\Delta\lambda_{1}=\delta\lambda_{1}=0$ (60)
$D\lambda_{2}=2(\lambda_{1}+\lambda_{2})(\sigma-\bar{\rho})$ (61)
$\Delta\lambda_{2}=0$ (62)
$\delta\lambda_{2}=2(\lambda_{1}+\lambda_{2})(\beta-\bar{\alpha})$ (63)
$D\lambda_{7}=2\lambda_{7}(\sigma-\bar{\rho})$ (64)
$\Delta\lambda_{7}=0$ (65)
$\delta\lambda_{7}=2\lambda_{7}(\beta-\bar{\alpha})$ (66)
Since we obtained the derivatives of $\lambda$’s we are able now to derive the integrability conditions of $K^{0}_{\mu\nu}$ for the first case. Based on the equation $Q+1=\lambda_{1}+\lambda_{2}+\lambda_{7}=0$ we can easily check that
$\frac{D(\lambda_{1}+\lambda_{2})}{\lambda_{1}+\lambda_{2}}=\frac{D\lambda_{7}}{\lambda_{7}}=2(\sigma-\bar{\rho})$ (67)
$\frac{\delta(\lambda_{1}+\lambda_{2})}{\lambda_{1}+\lambda_{2}}=\frac{\delta\lambda_{7}}{\lambda_{7}}=2(\beta-\bar{\alpha})$ (68)
5.1 Integrability conditions of $K^{0}_{\mu\nu}$
The integrability conditions for $K^{0}_{\mu\nu}$ form are given as follows.
$\Delta(\sigma-\bar{\rho})=(\alpha-\bar{\beta})(\pi+\bar{\pi}+\tau+\bar{\tau})$ (CR1:$\lambda_{7}$)
$2\delta(\sigma-\bar{\rho})+2D(\alpha-\bar{\beta})=(\sigma-\bar{\rho})(\bar{\alpha}+\beta-\bar{\pi})-(\alpha-\bar{\beta})(\sigma+\bar{\rho})$ (CR2:$\lambda_{7}$)
$\Delta(\alpha-\bar{\beta})=0$ (CR3:$\lambda_{7}$)
$(\delta-\bar{\delta})(\alpha-\bar{\beta})=0$ (CR4:$\lambda_{7}$)
The relations (55)-(59) and the integrability conditions provide us with further information crystallizes the usefulness of the entanglement of the Killing tensor. The latter along with the Newman-Penrose equations and the Bianchi identities form a solvable-overdetermined system of equations.
$D\rho-\bar{\delta}\kappa={\rho}(\rho+2\epsilon)+\sigma\bar{\sigma}-\kappa\left[2(\alpha+\bar{\beta})+(\alpha-\bar{\beta})-\pi+\tau\right]+\Phi_{00}$ (a)
$\delta\kappa-D\sigma=-(\rho+\bar{\rho}+2\epsilon)\sigma+\kappa\left[\tau-\bar{\pi}+2(\bar{\alpha}+\beta)-(\bar{\alpha}-\beta)\right]-\Psi_{o}$ (b)
$D\tau-\Delta\kappa=\rho(\tau+\bar{\pi})+\sigma(\pi+\bar{\tau})+\Psi_{1}$ (c)
$-\Delta\pi=\Psi_{3}$ (i)
$\bar{\delta}\pi=-\pi(\pi+\alpha-\bar{\beta})$ (g)
$\delta\tau-\Delta\sigma=\tau(\tau-\bar{\alpha}+\beta)$ (p)
$-\delta\pi=\pi(\bar{\pi}-\bar{\alpha}+\beta)+\Psi_{2}+2\Lambda$ (h)
$\Delta\rho-\bar{\delta}\tau=-\tau(\bar{\tau}+\alpha-\bar{\beta})-\Psi_{2}-2\Lambda$ (q)
$\delta\rho-\bar{\delta}\sigma=\rho(\bar{\alpha}+\beta)-\sigma(3\alpha-\bar{\beta})+\tau(\rho-\bar{\rho})-\Psi_{1}$ (k)
$0=\Psi_{3}$ (m)
$D\alpha-\bar{\delta}\epsilon=\alpha(\rho-\epsilon)+\beta\bar{\sigma}-\bar{\beta}\epsilon+\pi(\epsilon+\rho)$ (d)
$D\beta-\delta{\epsilon}=\sigma(\alpha+\pi)+\beta(\bar{\rho}-\bar{\epsilon})-\epsilon(\bar{\alpha}-\bar{\pi})+\Psi_{1}$ (e)
$\Delta\alpha=\Psi_{3}$ (r)
$\Delta\beta=0$ (o)
$\delta\alpha-\bar{\delta}\beta=\alpha(\bar{\alpha}-\beta)-\beta(\alpha-\bar{\beta})-\Psi_{2}+\Lambda$ (l)
$-\Delta\epsilon=\alpha(\tau+\bar{\pi})+\beta(\bar{\tau}+\pi)+\Psi_{2}-\Lambda+\tau\pi$ (f)
$0=\Psi_{4}$ (j)
The Bianchi identities with $\Psi_{3}=\Psi_{4}=0$ is given below.
$\bar{\delta}\Psi_{0}-D\Psi_{1}=(4\alpha-\pi)\Psi_{0}-2(2\rho+\epsilon)\Psi_{1}+3\kappa\Psi_{2}$ (I)
$\bar{\delta}\Psi_{1}-D\Psi_{2}=2(\alpha-\pi)\Psi_{1}-3\rho\Psi_{2}$ (II)
$\bar{\delta}\Psi_{2}=-3\pi\Psi_{2}$ (III)
$\Delta\Psi_{0}-\delta\Psi_{1}=-2(2\tau+\beta)\Psi_{1}+3\sigma\Psi_{2}$ (V)
$\Delta\Psi_{1}-\delta\Psi_{2}=-3\tau\Psi_{2}$ (VI)
$\Delta\Psi_{2}=0$ (VII)
5.2 A unique Petrov type D solution
At this point, we must ascertain that none of the remaining Weyl components are nullified. Now, we will prove that $\Psi_{2}$ is non-zero, thereby confirming that this solution is of Petrov type D. The real and imaginary part of $\Psi_{2}$ could be obtained by the comparison of NPE (p) with ($CR1:\lambda_{7}$) considering the relation $\epsilon=\rho-\bar{\sigma}$.
$\Psi_{2}+\bar{\Psi}_{2}-2\Lambda=(\alpha-\bar{\beta})(\pi+\bar{\pi}+\tau+\bar{\tau})+2\pi\bar{\pi}$ (69)
$\Psi_{2}-\bar{\Psi}_{2}=(\bar{\tau}-\tau)\left[2(\alpha-\bar{\beta})+\pi-\bar{\tau}\right]$ (70)
We are not be able to determine if this solution is a brand new type D solution, however the main characteristics of this type D solution is inferred in the downward relations.
$\kappa\sigma\rho\neq 0$ (71)
$\Psi_{4}=C_{\kappa\lambda\mu\nu}n^{\kappa}\bar{m}^{\lambda}n^{\mu}\bar{m}^{\nu}=0$ (72)
$\Psi_{3}=C_{\kappa\lambda\mu\nu}n^{\kappa}l^{\lambda}n^{\mu}\bar{m}^{\nu}=0$ (73)
Furthermore, we should note that any rotation around the null tetrad frame is inapplicable due to the non-preservation of the Killing tensor during rotation. Also, considering the aforementioned relations, there is a non-geodesic ($\kappa\neq 0$), a non-diverging ($\rho\neq 0$) null congruence characterized by shear ($\sigma\neq 0$) and our solution is of type D, with $\Psi_{2}\Psi_{0}\Psi_{1}\neq 0$, and the null vector $n^{\mu}$ serve as a Principal Null Direction (PND).
Finally, we performed the exact same derivations for $Q-1=\lambda_{1}+\lambda_{2}-\lambda_{7}=0$, and we arrived at the same Petrov type D, with the same characteristics of the solution to be hold.
Theorem 1.
A unique Petrov type D solution with a non-geodesic, shearing and diverging null congruence $n^{\mu}$ admits $K^{0}_{\mu\nu}$ canonical form of Killing tensor.
6 Solutions of $K^{1}_{\mu\nu}$ form
The $K^{1}_{\mu\nu}$ form has been handled in a unique manner. We nullify $\lambda_{7}$ and set $\lambda_{0}$ to be a constant equal to a value represented as $q=\pm 1$. This approach was chosen to achieve a similar form to that of Hauser-Malhiot (Paradigm) with the only difference being the constant $q$.
With these simplifications, we obtain a Jordan form of Killing tensor, as opposed to a diagonalized form with two double eigenvalues. Besides, it is widely known that the Jordan canonical form of a matrix embodies all the similar matrices of the family of matrices with the same eigenvalues except the “unique” member of this family, the diagonalized member of the family [Strang, 2022]. We shall proceed now with the Killing equation serves as a starting point,
$K^{1}_{\mu\nu}=qn_{\mu}n_{\nu}+\lambda_{1}(l_{\mu}n_{\nu}+n_{\mu}l_{\nu})+\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu})$ (74)
To acquire solutions of NPE, first, we have to consider additional mathematical conditions. In our case these represent the integrability conditions of the eigenvalues $\lambda_{1}$, $\lambda_{2}$ of our Killing tensor which come by the Killing equation.
$K_{(\mu\nu;\alpha)}=0$ (75)
The Killing equation provides us with the following relations. They will be used along with the CR in order to give us the IC of the eigenvalues $\lambda_{1},\lambda_{2}$,
$\nu=\sigma=\lambda=0$ (76)
$q(\epsilon+\bar{\epsilon})=0\hskip 28.45274ptq\neq 0$ (77)
$(\lambda_{1}+\lambda_{2})\kappa=q(\bar{\alpha}+\beta+\bar{\pi})$ (78)
The directional derivatives for eigenvalues $\lambda_{1}$, $\lambda_{2}$ turned out to be the following
$D\lambda_{1}=q(\gamma+\bar{\gamma})$ (79)
$\Delta\lambda_{1}=0$ (80)
$\delta\lambda_{1}=(\lambda_{1}+\lambda_{2})(\bar{\pi}-\tau)$ (81)
$D\lambda_{2}=q(\mu+\bar{\mu})-(\lambda_{1}+\lambda_{2})(\rho+\bar{\rho})$ (82)
$\Delta\lambda_{2}=(\lambda_{1}+\lambda_{2})(\mu+\bar{\mu})$ (83)
$\delta\lambda_{2}=0$ (84)
The relation (78) could be used defining a factor $Q$ with its directional derivatives
$Q\equiv\frac{q}{\lambda_{1}+\lambda_{2}}=\frac{\kappa}{\bar{\alpha}+\beta+\bar{\pi}}$ (85)
$DQ=Q[\rho+\bar{\rho}-Q(\gamma+\bar{\gamma}+\mu+\bar{\mu})]$ (86)
$\Delta Q=-Q(\mu+\bar{\mu})$ (87)
$\delta Q=-Q(\bar{\pi}-\tau)$ (88)
The factor $Q$ was proved helpful in the treatment of the IC and it is a real function since it depends only from real eigenvalues.
6.1 Integrabillity Conditions of $K^{1}_{\mu\nu}$ form
IC come to surface by acting of the CR upon to the eigenvalues. Additionally, CR is the Lie bracket of the basis vectors. We choose to separate the IC in two parts using the factor $Q$.
Integrability Conditions of Eigenvalue $\lambda_{1}$
$Q[\delta(\gamma+\bar{\gamma})-(\gamma+\bar{\gamma})(\bar{\alpha}+\beta-\tau)-(\mu+\bar{\mu})(\bar{\pi}-\tau)]=D(\bar{\pi}-\tau)-(\bar{\pi}-\tau)(\rho+\bar{\rho})-(\bar{\pi}-\tau)(2\epsilon+\bar{\rho})$ ($CR1:\lambda_{1}$)
$Q[\Delta(\gamma+\bar{\gamma})-(\gamma+\bar{\gamma})^{2}]=2(\tau\bar{\tau}-\pi\bar{\pi})$ ($CR2:\lambda_{1}$)
$\Delta(\bar{\pi}-\tau)-(\bar{\pi}-\tau)[(\gamma-\bar{\gamma})-(2\mu+\bar{\mu})]=0$ ($CR3:\lambda_{1}$)
$\delta(\pi-\bar{\tau})-\bar{\delta}(\bar{\pi}-\tau)+(\bar{\pi}-\tau)(\alpha-\bar{\beta})-(\pi-\bar{\tau})(\bar{\alpha}-\beta)=Q(\gamma+\bar{\gamma})(\mu-\bar{\mu})$ ($CR4:\lambda_{1}$)
Integrability Conditions of Eigenvalue $\lambda_{2}$
$Q[\delta(\mu+\bar{\mu})-2(\mu+\bar{\mu})(\bar{\alpha}+\beta)]=\delta(\rho+\bar{\rho})-(\rho+\bar{\rho})[(\bar{\alpha}+\beta-\bar{\pi})-(\bar{\pi}-\tau)]$ ($CR1:\lambda_{2}$)
$Q[\Delta(\mu+\bar{\mu})-(\mu+\bar{\mu})[(\mu+\bar{\mu})+2(\gamma+\bar{\gamma})]]=D(\mu+\bar{\mu})+\Delta(\rho+\bar{\rho})-(\rho+\bar{\rho})(\gamma+\bar{\gamma})$ ($CR2:\lambda_{2}$)
$\delta(\mu+\bar{\mu})+(\mu+\bar{\mu})[(\bar{\alpha}+\beta-\tau)+(\bar{\pi}-\tau)]=0$ ($CR3:\lambda_{2}$)
$2(\mu\bar{\rho}-\bar{\mu}\rho)=Q(\mu+\bar{\mu})(\mu-\bar{\mu})$ ($CR4:\lambda_{2}$)
6.2 Classes of solutions
Consider now the $K^{1}$ Killing form.
$K^{1}=q\tilde{\theta}^{1}\otimes\tilde{\theta}^{1}+\lambda_{1}(\tilde{\theta}^{1}\otimes\tilde{\theta}^{2}+\tilde{\theta}^{2}\otimes\tilde{\theta}^{1})+\lambda_{2}(\tilde{\theta}^{3}\otimes\tilde{\theta}^{4}+\tilde{\theta}^{4}\otimes\tilde{\theta}^{3});\hskip 14.22636ptq=\pm 1$
This Killing form dictates that the only non-zero rotation parameter is $t=ib$. With the usage of the general key relations (i)-(iv) of chapter 4 we get the following relations.
$\Psi_{2}-\Lambda=\tau\pi$ (i)
$\Psi_{1}=\kappa\mu$ (ii)
$\Psi_{2}-\Lambda=\mu\rho$ (iii)
$\mu\tau=0$ (iv)
The rotation provides us with these useful relations which connect the spin coefficients along with the Weyl components. This result depends mainly on the form of the Killing Tensor since we demand by the rotation to preserve the Killing tensor invariant. Essentially, the annihilation of $\lambda_{7}$ allows us to obtain simplifications such that.
Thus, the key relations could help us to obtain the Petrov types of this solution initiating by the key relation (iv). The possible classes are three $\mu=0\neq\tau$, $\mu=0=\tau$ and $\mu\neq 0=\tau$.
6.2.1 Class I: $\mu=0\neq\tau$
The annihilation of $\mu$ has an immediate impact at key relations in the first place. Beginning with the condition $\tau\neq 0$, the first relation yields the annihilation of $\pi$ since $\Psi_{2}=\Lambda$ due to relation (iii),
$0=\pi$ (i)
$\Psi_{1}=0$ (ii)
$\Psi_{2}-\Lambda=0$ (iii)
$\mu=0\neq\tau$ (iv)
Afterwards, we will try to plug these into Newman-Penrose equations and into the integrability conditions. We take the annihilation of every Weyl component except $\Psi_{0}$. We already know that $\nu=0$ nullifies $\Psi_{4}$. Next, the simultaneous nullification of $\pi$ and $\mu$ along with the NPE (m) and (h) clarifies that the solution is of type N since the only survivor is the $\Psi_{0}$.
6.2.2 Class II: $\mu=0=\tau$
The second class is determined by the following relations
$0=\tau$ (i)
$\Psi_{1}=0$ (ii)
$\Psi_{2}-\Lambda=0$ (iii)
$\mu=0=\tau$ (iv)
This combination does not make any difference. Again, we have nullification of $\Psi_{4}$ due to NPE (j). Additionally, the IC $(CR3:\lambda_{1})$ along with NPE (i) nullifies $\Psi_{3}$ and the ($CR2:\lambda_{2}$) along with NPE (q) nullifies $\Psi_{2}$. This case is also characterized as Type N.
6.2.3 Class III: $\mu\neq 0=\tau$
The key relations for the last class follow
$\Psi_{2}-\Lambda=0$ (i)
$\Psi_{1}=\kappa\mu$ (ii)
$0=\rho$ (iii)
$\mu\neq 0=\tau$ (iv)
This case is a little trickier but it is proved to have the same Petrov type. The annihilation of $\nu$ gives $\Psi_{4}$. The IC $(CR4:\lambda_{2})$ yields two possibilities,
$(\mu+\bar{\mu})(\mu-\bar{\mu})=0$ (89)
The only option is $\mu-\bar{\mu}=0$ since the simultaneous annihilation of $\rho$ and $\mu+\bar{\mu}$ gives $\lambda_{2}=const$ resulting in a canonical form without two double eigenvalues. The latter along with NPE (k) annihilate $\Psi_{1}$. Subsequently, the NPE (q) nullifies the $\Psi_{2}$ eventually, so the only non-zero Weyl components are $\Psi_{0}$ and $\Psi_{3}$. Initiating by NPE (i) and IC $(CR3:\lambda_{1})$ we get
$\Psi_{3}=2\pi\mu.$ (90)
By $(CR1:\lambda_{2})$ and $(CR3:\lambda_{2})$ we take
$\pi=\alpha+\bar{\beta}$ (91)
In this point we introduce the Bianchi identity (II) which yields with two possible outcomes,
$\kappa\Psi_{3}=0$ (92)
It is obvious that the annihilation of $\Psi_{3}$ gives the bespoken result. The Petrov type is determined as type N. But the other option, where $\kappa=Q(\bar{\alpha}+\beta+\bar{\pi})=0$ along with (90) and (91) gives $\pi=\alpha+\bar{\beta}=0$. Then, the space is characterized as conformally flat since the last relation annihilates $\Psi_{3}$ and $\kappa=0$ implying that $\Psi_{0}=0$ (NPE (b)).
Table 1: Rotation around $\mu\tau=0$
Type N Type N Type N
$\mu=0\neq\tau$ $\mu=0=\tau$ $\mu\neq 0=\tau$
$\Psi_{0}\neq 0$ $\Psi_{0}\neq 0$ $\Psi_{0}\neq 0$
$d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$
At last, the implication of the rotation provides us with simplifications paving the way to the following theorem.
Theorem 2.
Petrov type N solution with a shear-free (non-diverging in one case) null congruence admits $K^{1}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{7}=0$.
7 Solutions of $K^{2}_{\mu\nu}$and $K^{3}_{\mu\nu}$ form
Our objective is to determine the geometrical characterization (Petrov types) of the gravitational fields associated with the derived solutions, assuming the existence of $K^{2}$ and $K^{3}$ Killing forms with $\lambda_{7}=0$. In the next chapter we will exhibit the case where $\lambda_{0}=0$.
$K_{(\mu\nu;\alpha)}=0$ (93)
Defining the factor $q=\pm 1$, we consider a unified approach for both canonical forms $K^{2}$ and $K^{3}$. The only difference in the $K^{2}$ and $K^{3}$ forms is the $-1$ in the $K_{22}$ component. Obviously, we get the $K^{2}_{\mu\nu}$ for $q=+1$ and the $K^{3}_{\mu\nu}$ for $q=-1$.
$K^{2,3}_{\mu\nu}=\lambda_{0}(n_{\mu}n_{\nu}+ql_{\mu}l_{\nu})+\lambda_{1}(l_{\mu}n_{\nu}+n_{\mu}l_{\nu})-\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu})$ (94)
This modification allows us to study the two forms simultaneously. The Killing equation of the Killing form yields the annihilation of $\sigma$ and $\lambda$ and the directional derivatives of $\lambda_{0}$, $\lambda_{1}$, $\lambda_{2}$.
$\sigma=\lambda=0$ (95)
$D\lambda_{0}=2\lambda_{0}(\epsilon+\bar{\epsilon})$ (96)
$\Delta\lambda_{0}=-2\lambda_{0}(\gamma+\bar{\gamma})$ (97)
$\delta\lambda_{0}=2\left[\lambda_{0}(\bar{\alpha}+\beta+\bar{\pi})-\kappa(\lambda_{1}+\lambda_{2})\right]$ (98)
$\delta\lambda_{0}=2\left[-\lambda_{0}(\bar{\alpha}+\beta+\tau)+q\bar{\nu}(\lambda_{1}+\lambda_{2})\right]$ (99)
$\delta\lambda_{0}=\lambda_{0}(\bar{\pi}-\tau)-(\kappa-q\bar{\nu})(\lambda_{1}+\lambda_{2})$ (100)
$D\lambda_{1}=2\lambda_{0}(\gamma+\bar{\gamma})$ (101)
$\Delta\lambda_{1}=-2q\lambda_{0}(\epsilon+\bar{\epsilon})$ (102)
$\delta\lambda_{1}=-q\lambda_{0}(\kappa-q\bar{\nu})+(\lambda_{1}+\lambda_{2})(\bar{\pi}-\tau)$ (103)
$D\lambda_{2}=\lambda_{0}(\mu+\bar{\mu})-(\lambda_{1}+\lambda_{2})(\rho+\bar{\rho})$ (104)
$\Delta\lambda_{2}=-q\lambda_{0}(\rho+\bar{\rho})-(\lambda_{1}+\lambda_{2})(\mu+\bar{\mu})$ (105)
$\delta\lambda_{2}=0$ (106)
Relations (98) and (99) allow us to define the factor $Q$.
$Q\equiv\frac{\lambda_{0}}{\lambda_{1}+\lambda_{2}}=\frac{\kappa+q\bar{\nu}}{2(\bar{\alpha}+\beta)+\bar{\pi}+\tau}$ (107)
$DQ=Q(2(\epsilon+\bar{\epsilon})+(\rho+\bar{\rho}))-Q^{2}(2(\gamma+\bar{\gamma})+(\mu+\bar{\mu}))$ (108)
$\Delta Q=-Q(2(\gamma+\bar{\gamma})+(\mu+\bar{\mu}))+qQ^{2}(2(\epsilon+\bar{\epsilon})+(\rho+\bar{\rho}))$ (109)
$\delta Q=(qQ^{2}-1)(\kappa-q\bar{\nu})$ (110)
The factor $Q$ is proved helpful during the treatment of the IC and it is a real scalar function since it depends solely on real scalars.
7.1 Integrabillity Conditions of $K^{2}_{\mu\nu}$ and $K^{3}_{\mu\nu}$ forms
We use the commutators of the tetrads in order to obtain the integrability conditions of Killing tensor. As we mentioned in subsection 2.2, the commutation relations are equivalent with the Lie bracket of the null tetrads. We choose to separate the integrability conditions using the factor $Q$.
Integrability Conditions of $\lambda_{0}$
$2Q[D(\gamma+\bar{\gamma})+\Delta(\epsilon+\bar{\epsilon})+\pi\bar{\pi}-\tau\bar{\tau}]=-[(\pi+\bar{\tau})(q\bar{\nu}-\kappa)+(\bar{\pi}+\tau)(q\nu-\bar{\kappa})]$ ($CR1:\lambda_{0}$)
$Q[2[\delta(\epsilon+\bar{\epsilon})-(\epsilon+\bar{\epsilon})(\bar{\alpha}+\beta-\bar{\pi})]-[D(\bar{\pi}-\tau)-(\bar{\pi}-\tau)(\bar{\rho}+\epsilon-\bar{\epsilon})]+2\kappa(\gamma+\bar{\gamma})-(q\bar{\nu}-\kappa)[2(\gamma+\bar{\gamma})\\ +(\mu+\bar{\mu})]]=D(q\bar{\nu}-\kappa)-(q\bar{\nu}-\kappa)[2\epsilon+\bar{\rho}+\epsilon+\bar{\epsilon}+\rho+\bar{\rho}]$ (111)
$Q[2[\delta(\gamma+\bar{\gamma})+(\gamma+\bar{\gamma})(\bar{\alpha}+\beta-\tau)]+[\Delta(\bar{\pi}-\tau)+(\bar{\pi}-\tau)(\mu-\gamma+\bar{\gamma})]-2\bar{\nu}(\epsilon+\bar{\epsilon})-q(q\bar{\nu}-\kappa)[2(\epsilon+\bar{\epsilon})\\ +\rho+\bar{\rho}]]=\Delta(\kappa-q\bar{\nu})+(\kappa-q\bar{\nu})[2(\gamma+\bar{\gamma})+(\mu+\bar{\mu})+\mu-\gamma+\bar{\gamma}]$ (112)
$Q[\bar{\delta}(\bar{\pi}-\tau)-\delta(\pi-\bar{\tau})-(\bar{\pi}-\tau)(\alpha-\bar{\beta})+(\pi-\bar{\tau})(\bar{\alpha}-\beta)+2[(\epsilon+\bar{\epsilon})(\mu-\bar{\mu})-(\gamma+\bar{\gamma})(\rho-\bar{\rho})]]\\ =\delta(q\nu-\bar{\kappa})-\bar{\delta}(q\bar{\nu}-\kappa)+(q\bar{\nu}-\kappa)(\alpha-\bar{\beta})-(q\nu-\bar{\kappa})(\bar{\alpha}-\beta)$ (113)
Integrability Conditions of $\lambda_{1}$
$Q[\Delta(\gamma+\bar{\gamma})-3(\gamma+\bar{\gamma})^{2}+q[D(\epsilon+\bar{\epsilon})+3(\epsilon+\bar{\epsilon})^{2}]+\frac{q}{2}[(\pi+\bar{\tau})(q\bar{\nu}-\kappa)+(\bar{\pi}+\tau)(q\nu-\bar{\kappa})]]=-(\pi\bar{\pi}-\tau\bar{\tau})$ ($CR1:\lambda_{1}$)
$Q[2[\delta(\gamma+\bar{\gamma})-(\gamma+\bar{\gamma})(\bar{\alpha}+\beta-\bar{\pi})]-q[D(q\bar{\nu}-\kappa)+(q\bar{\nu}-\kappa)(\epsilon+3\bar{\epsilon}+\bar{\rho})-2\kappa(\epsilon+\bar{\epsilon})]]\\ =D(\bar{\pi}-\tau)-(\bar{\pi}-\tau)(\rho+2\bar{\rho}+\epsilon-\bar{\epsilon})-2(\gamma+\bar{\gamma})(q\bar{\nu}-\kappa)$ (114)
$Q[2q[\delta(\epsilon+\bar{\epsilon})+(\epsilon+\bar{\epsilon})(\bar{\alpha}+\beta-\tau)]+q[\Delta(q\bar{\nu}-\kappa)-(q\bar{\nu}-\kappa)(3\gamma+\bar{\gamma}-\mu)]-2\bar{\nu}(\gamma+\bar{\gamma})]\\ =-[\Delta(\bar{\pi}-\tau)+(\bar{\pi}-\tau)(2\mu+\bar{\mu}-\gamma+\bar{\gamma})+2q(q\bar{\nu}-\kappa)(\epsilon+\bar{\epsilon})]$ (115)
$Q[q[\delta(q\nu-\bar{\kappa})-\bar{\delta}(q\bar{\nu}-\kappa)+(q\bar{\nu}-\kappa)(\alpha-\bar{\beta})-(q\nu-\bar{\kappa})(\bar{\alpha}-\beta)]+2[q(\epsilon+\bar{\epsilon})(\rho-\bar{\rho})-(\gamma+\bar{\gamma})(\mu-\bar{\mu})]\\ =\bar{\delta}(\bar{\pi}-\tau)-\delta(\pi-\bar{\tau})-(\bar{\pi}-\tau)(\alpha-\bar{\beta})+(\pi-\bar{\tau})(\bar{\alpha}-\beta)$ (116)
Integrability Conditions of $\lambda_{2}$
$Q[[\Delta(\mu+\bar{\mu})-(\mu+\bar{\mu})-5(\gamma+\bar{\gamma})]+q[D(\rho+\bar{\rho})+(\rho+\bar{\rho})[(\rho+\bar{\rho})-5(\epsilon+\bar{\epsilon})]]]\\ =\Delta(\rho+\bar{\rho})-(\rho+\bar{\rho})(\gamma+\bar{\gamma})+D(\mu+\bar{\mu})+(\mu+\bar{\mu})(\epsilon+\bar{\epsilon})$ (117)
$Q[\delta(\mu+\bar{\mu})-(\mu+\bar{\mu})[(\bar{\alpha}+\beta+\tau)-2\bar{\pi}]+q(\rho+\bar{\rho})(2\kappa-q\bar{\nu})]\\ =\delta(\rho+\bar{\rho})-(\rho+\bar{\rho})[\bar{\alpha}+\beta+\tau-2\bar{\pi}]+(\mu+\bar{\mu})(2\kappa-q\bar{\nu})$ (118)
$qQ[\delta(\rho+\bar{\rho})+(\rho+\bar{\rho})[\bar{\alpha}+\beta+\bar{\pi}-2\tau]+(\mu+\bar{\mu})(\kappa-2q\bar{\nu})]\\ =\delta(\mu+\bar{\mu})+(\mu+\bar{\mu})[\bar{\alpha}+\beta+\bar{\pi}-2\tau]+q(\rho+\bar{\rho})(\kappa-2q\bar{\nu})$ (119)
$Q[(\mu+\bar{\mu})(\mu-\bar{\mu})-q(\rho+\bar{\rho})(\rho-\bar{\rho})]=(\mu-\bar{\mu})(\rho+\bar{\rho})-(\rho-\bar{\rho})(\mu+\bar{\mu})$ (120)
7.2 Classes of solutions
We choose to take advantage of the conformal symmetry of a rotation around one of the real null vectors (Chapter 3). Subsequently, we require that the Killing tensor remains invariant under rotation, as discussed.
$K=\lambda_{0}(\tilde{\theta}^{1}\otimes\tilde{\theta}^{1}+q\tilde{\theta}^{2}\otimes\tilde{\theta}^{2})+\lambda_{1}(\tilde{\theta}^{1}\otimes\tilde{\theta}^{2}+\tilde{\theta}^{2}\otimes\tilde{\theta}^{1})+\lambda_{2}(\tilde{\theta}^{3}\otimes\tilde{\theta}^{4}+\tilde{\theta}^{4}\otimes\tilde{\theta}^{3});\hskip 14.22636ptq=\pm 1$
It is easy for someone to prove that the only non-zero rotation parameter is $t=ib$. As a matter of fact, we obtain the exact same key relations since the null rotation is applied with the exact same way as in the previous chapter.
Moving forward, we shall present the classes of solutions are classified by the usage of key relations. Initiating by $\mu\tau=0$ the NPE, IC, BI, provide us with three main classes of solutions. Also, we aim to study solutions in which only one of the components of Killing tensor, is allowed to be constant. It would be helpful for the reader if at this point we array the most useful NPE and the $(CR4:\lambda_{2})$.
$\delta\kappa=\kappa(\tau-\bar{\pi}+2(\bar{\alpha}+\beta)-(\bar{\alpha}-\beta))-\Psi_{o}$ (b)
$\bar{\delta}\pi=-\pi(\pi+\alpha-\bar{\beta})+\nu\bar{\kappa}$ (g)
$\delta\tau=\tau(\tau-\bar{\alpha}+\beta)-\bar{\nu}\kappa$ (p)
$\delta\rho=\rho(\bar{\alpha}+\beta)+\tau(\rho-\bar{\rho})+\kappa(\mu-\bar{\mu})-\Psi_{1}$ (k)
$\bar{\delta}\mu=-\mu(\alpha+\bar{\beta})-\pi(\mu-\bar{\mu})-\nu(\rho-\bar{\rho})+\Psi_{3}$ (m)
$\bar{\delta}\nu=-\nu(2(\alpha+\bar{\beta})+(\alpha-\bar{\beta})+\pi-\bar{\tau})+\Psi_{4}$ (j)
$Q[(\mu+\bar{\mu})(\mu-\bar{\mu})-q(\rho+\bar{\rho})(\rho-\bar{\rho})]=(\mu-\bar{\mu})(\rho+\bar{\rho})-(\rho-\bar{\rho})(\mu+\bar{\mu})$ (121)
7.2.1 Class I : $\mu=0$
The annihilation of $\mu$ gives the first class of solution. Considering that $\mu=0$ the relation $(CR4:\lambda_{2})$ gives $(\rho+\bar{\rho})(\rho-\bar{\rho})=0$. According to IC (104)-(106) the derivative of the eigenvalue $\lambda_{2}$ depends only from the real parts of $\mu$ and $\rho$, so our priority is to avoid $\rho+\bar{\rho}=0$. Unavoidably, the BI (II) implies that the derivative of $\lambda_{2}$ vanishes. However, from NPE (m) we take $\Psi_{3}=\nu(\rho-\bar{\rho})=0$.
In the following Tables 2, 3 every column represents different solutions according to different choices which are ordered by the key relations and the BI (II). The columns of the tables contain the main characteristics of our solutions. The second and third column of Table 2 is distinguished by the different choices that take place due to the BI (III) and BI (VI).
Table 2: $\rho-\bar{\rho}=0=\rho+\bar{\rho}$
Type D Type N Type N
$\rho=0\neq\Psi_{2}$ $\rho+\bar{\rho}=0=\Psi_{2}$ $\rho+\bar{\rho}=0=\Psi_{2}$
$\kappa\nu=\pi\tau$ $\nu=0=\pi\tau$ $\kappa=0=\pi\tau$
$\Psi_{0}\Psi_{4}={9\Psi_{2}}^{2}$ $\Psi_{0}\neq 0$ $\Psi_{4}\neq 0$
$\Psi_{1}=\Psi_{3}=0$ $\Psi_{1}=\Psi_{2}=\Psi_{3}=\Psi_{4}=0$ $\Psi_{0}=\Psi_{1}=\Psi_{2}=\Psi_{3}=0$
$d\lambda_{2}=0$ $d\lambda_{2}=0$ $d\lambda_{2}=0$
The other choice where $\rho+\bar{\rho}\neq 0=\Psi_{2}$ yields type N solutions only where the eigenvalue $\lambda_{2}$ is not a constant. The combination of BI (III) with BI (VI) determines the non-zero Weyl component.
Table 3: $\rho-\bar{\rho}=0\neq\rho+\bar{\rho}$
Type N Type N Type N Type N Type N Type N
$\nu=0=\tau$ $\nu=0=\pi$ $\kappa=0=\tau$ $\kappa=0=\pi$ $\nu=\pi=\tau=0$ $\kappa=\pi=\tau=0$
$\Psi_{0}\neq 0$ $\Psi_{4}\neq 0$ $\Psi_{0}\neq 0$ $\Psi_{4}\neq 0$ $\Psi_{0}\neq 0$ $\Psi_{4}\neq 0$
$d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}\neq 0$
7.2.2 Class II: $\mu=0=\tau$
As before, the NPE (m) yields $\Psi_{3}=\nu(\rho-\bar{\rho})=0$. Considering the key relation (ii) we obtain $\Psi_{1}=0$, similarly considering NPE (i) we obtain $\kappa\nu=0$. Thus, it is obvious that the Class II consists a subset of Class I. These solutions are the same type N solutions of the previous class with the only difference to be that $\tau=0$.
7.2.3 Class III: $\tau=0$
In this class we encountered new algebraically special solutions. NPE (p) for $\tau=0$ yields $\bar{\kappa}\nu=0$. The branch where $\mu=0$ is already known from the other two classes where concerns type N solutions with $\tau=0$. On the other hand, the case of $\mu\neq 0=\rho$ yields two Type III solutions.
The IC $(CR4:\lambda_{2})$ plays a crucial role in this scene since the annihilation of $\rho$ implies $(\mu+\bar{\mu})(\mu-\bar{\mu})=0$. Although, the constraint $\mu\neq 0$ leads us to two separate solutions for case $\kappa=0\neq\nu$, which are both of Type III. The other case, where both $\kappa$ and $\nu$ are zero, concerns two solutions where only $\Psi_{3}$ is not equal to zero determining that the type of the solutions are also of Type III.
The last branch of solutions contains the case $\kappa\neq 0=\nu$. It should be noted that NPE (k) with $\rho=0$ determines $\Psi_{1}$, which annihilates the real part of $\mu$ which is equivalent with $d\lambda_{2}=0$. Then NPE (k) is given by the following relation.
$\Psi_{1}=\kappa(\mu-\bar{\mu})$ (k)
Table 4: $\mu\neq 0=\rho$
$\mu-\bar{\mu}=0$ $\mu+\bar{\mu}=0$ $\mu-\bar{\mu}=0$ $\mu+\bar{\mu}=0$
Type III Type III Type III Type III
$\kappa=0\neq\nu$ $\kappa=0\neq\nu$ $\kappa=0=\nu$ $\kappa=0=\nu$
$\Psi_{3}\neq 0\neq\Psi_{4}$ $\Psi_{3}\neq 0\neq\Psi_{4}$ $\Psi_{3}\neq 0=\Psi_{4}$ $\Psi_{3}\neq 0=\Psi_{4}$
$d\lambda_{2}\neq 0$ $d\lambda_{2}=0$ $d\lambda_{2}\neq 0$ $d\lambda_{2}=0$
The Class III is presented in Table 4. All these cases are characterized by $\Psi_{2}=\Lambda=0$ which arises from NPE (q) for $\rho=\tau=0$, while in these type III solutions $\Psi_{1}=0$. Finally, we postulate the following two theorems.
Theorem 3.
Petrov types III, N solutions with a shear-free, diverging (and in some cases geodesic) null congruence admit both $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$.
Theorem 4.
Petrov type D solution with a shear-free, diverging and non-geodesic null congruence admit both $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$.
8 Solutions of $K^{4}_{\mu\nu}$ form
In line with the previous chapters we study one more form which does not belong to the four canonical forms of Killing tensor although is a subcase of $K^{1}_{\mu\nu},K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ forms and it is obtained when $\lambda_{0}$ equals to zero. We wanted to investigate the virtue of the solutions that one could obtain using the other possible reduction of Killing forms $K^{1},K^{2},K^{3}$, namely, when $\lambda_{0}=0$.
With reference to Remark 3 we are able to obtain somehow different simplifications due to the rotation around the null tetrad frame. We initiate our analysis defining the factor $Q$ in the same fashion as in Chapter 5.
$Q\equiv\frac{\lambda_{7}}{\lambda_{1}+\lambda_{2}}$ (122)
The Killing form is represented in terms of the null tetrads as follows.
$K^{4}_{\mu\nu}=\lambda_{1}(n_{\mu}l_{\nu}+l_{\mu}n_{\nu})+\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu})+\lambda_{7}(m_{\mu}m_{\nu}+\bar{m}_{\mu}\bar{m}_{\nu})$ (123)
The Killing equation yields
$\kappa(\lambda_{1}+\lambda_{2})+\bar{\kappa}\lambda_{7}=0$ (124)
$\nu(\lambda_{1}+\lambda_{2})+\bar{\nu}\lambda_{7}=0$ (125)
$D\lambda_{1}=\Delta\lambda_{1}=0$ (126)
$\delta\lambda_{1}=(\bar{\pi}-\tau)(\lambda_{1}+\lambda_{2})+(\pi-\bar{\tau})\lambda_{7}$ (127)
$D\lambda_{2}=-(\rho+\bar{\rho})(\lambda_{1}+\lambda_{2}-(\sigma+\bar{\sigma})\lambda_{7}$ (128)
$\Delta\lambda_{2}=(\mu+\bar{\mu})(\lambda_{1}+\lambda_{2})+(\lambda+\bar{\lambda})\lambda_{7}$ (129)
$\delta\lambda_{2}=2\lambda_{7}(\alpha-\bar{\beta})$ (130)
$D\lambda_{7}=-2\bar{\sigma}(\lambda_{1}+\lambda_{2})-2\rho\lambda_{7}$ (131)
$\Delta\lambda_{7}=2\lambda(\lambda_{1}+\lambda_{2})+2\bar{\mu}\lambda_{7}$ (132)
$\delta\lambda_{7}=-2\lambda_{7}(\bar{\alpha}-\beta)$ (133)
8.1 Integrability Conditions of $K^{4}_{\mu\nu}$ form
Integrability Conditions of $\lambda_{1}$
$02(\pi\bar{\pi}-\tau\bar{\tau})+Q\left[(\pi+\bar{\tau})(\pi-\bar{\tau})+(\bar{\pi}+\tau)(\bar{\pi}-\tau)\right]=$ 0 ($CR1:\lambda_{1}$)
$D(\bar{\pi}-\tau)-(\bar{\pi}-\tau)(\rho+2\bar{\rho})-(\pi-\bar{\tau})(\sigma+2\bar{\sigma})\\ +Q\left[D(\pi-\bar{\tau})-(\bar{\pi}-\tau)(2\sigma+\bar{\sigma})-(\pi-\bar{\tau})(2\rho+\bar{\rho})\right]=0$ (134)
$\Delta(\bar{\pi}-\tau)+(\bar{\pi}-\tau)(2\mu+\bar{\mu})+(\pi-\bar{\tau})(2\lambda+\bar{\lambda})\\ +Q\left[\Delta(\pi-\bar{\tau})+(\bar{\pi}-\tau)(\lambda+2\bar{\lambda})+(\pi-\bar{\tau})(\mu+2\bar{\mu})\right]=0$ (135)
$\bar{\delta}(\bar{\pi}-\tau)-\delta(\pi-\bar{\tau})+Q\left[\bar{\delta}(\pi-\bar{\tau})-\delta(\bar{\pi}-\tau)+(\bar{\pi}-\tau)^{2}-(\pi-\bar{\tau})^{2}\right]=0$ ($CR4:\lambda_{1}$)
Integrability Conditions of $\lambda_{2}$
$\Delta(\rho+\bar{\rho})+D(\mu+\bar{\mu})-(\gamma+\bar{\gamma})(\rho+\bar{\rho})+(\epsilon+\bar{\epsilon})(\mu+\bar{\mu})+2(\lambda\sigma-\bar{\lambda}\bar{\sigma})\\ +Q\left[\Delta(\sigma+\bar{\sigma})+D(\lambda+\bar{\lambda})+(\lambda+\bar{\lambda})(\epsilon+\bar{\epsilon}-\rho+\bar{\rho})-(\sigma+\bar{\sigma})(\gamma+\bar{\gamma}+\mu-\bar{\mu})\right]=0$ (136)
$\delta(\rho+\bar{\rho})+(\rho+\bar{\rho})(2\bar{\pi}-\tau-\bar{\alpha}-\beta)+\kappa(\mu+\bar{\mu})\\ +Q\left[\delta(\sigma+\bar{\sigma})+(\pi-\bar{\tau})(\rho+\bar{\rho})-(\sigma+\bar{\sigma})(\bar{\alpha}+\beta-\bar{\pi})+\kappa(\lambda+\bar{\lambda})\right]=0$ (137)
$\delta(\mu+\bar{\mu})+(\mu+\bar{\mu})(\bar{\pi}-2\tau+\bar{\alpha}+\beta)-\bar{\nu}(\rho+\bar{\rho})\\ Q\left[\delta(\lambda+\bar{\lambda})+(\mu+\bar{\mu})(\bar{\pi}-\tau)-(\lambda+\bar{\lambda})(\tau-\bar{\alpha}-\beta)-\bar{\nu}(\sigma+\bar{\sigma})\right]=0$ (138)
$(\mu-\bar{\mu})(\rho+\bar{\rho})-(\rho-\bar{\rho})(\mu+\bar{\mu})+Q\left[(\mu-\bar{\mu})(\sigma+\bar{\sigma})-(\rho-\bar{\rho})(\lambda+\bar{\lambda})\right]=0$ ($CR4:\lambda_{2}$)
Integrability Conditions of $\lambda_{7}$
$\Delta\bar{\sigma}+D\lambda-\bar{\sigma}(\gamma+\bar{\gamma}-\mu+\bar{\mu})+\lambda(\epsilon+\bar{\epsilon}+\rho-\bar{\rho})+Q\left[D\bar{\mu}+\Delta\rho-(\lambda\sigma-\bar{\lambda}\bar{\sigma})+\bar{\mu}(\epsilon+\bar{\epsilon})-\rho(\gamma+\bar{\gamma})\right]=0$ ($CR1:\lambda_{7}$)
$\delta\bar{\sigma}+\bar{\sigma}(2\bar{\pi}-\tau-\bar{\alpha}-\beta)+\kappa\lambda+Q\left[\delta\rho+\bar{\sigma}(\pi-\bar{\tau})-\rho(\bar{\alpha}+\beta-\bar{\pi})+\kappa\bar{\mu}\right]=0$ ($CR2:\lambda_{7}$)
$\delta\lambda+\lambda(\bar{\pi}-2\tau+\bar{\alpha}+\beta)-\bar{\sigma}\bar{\nu}+Q\left[\delta\bar{\mu}+\lambda(\pi-\bar{\tau})-\rho\bar{\nu}-\bar{\mu}(\tau-\bar{\alpha}-\beta)\right]=0$ ($CR3:\lambda_{7}$)
$\bar{\sigma}(\mu-\bar{\mu})-\lambda(\rho-\bar{\rho})+Q\left[\rho(\mu-\bar{\mu})-\bar{\mu}(\rho-\bar{\rho})\right]=0$ ($CR4:\lambda_{7}$)
8.2 Classes of solutions
Similarly, we choose $l^{\mu}$ to be fixed during the rotation. Afterward, we demand that the Killing tensor remains invariant under the rotation but in this case the non-zero rotation parameter is proved to be $t=a$ instead of $t=ib$. The latter provides us with different simplifications due to the annihilation of the tilded spin coefficients. Namely, we get the key relations which unfold the branches of the solutions,
$\Psi_{2}-\Lambda=\kappa\nu-\tau\pi$ (i)
$\Psi_{1}=\kappa\mu-\sigma\pi$ (ii)
$\Psi_{2}-\Lambda=\mu\rho-\sigma\lambda$ (iii)
$\mu\tau-\sigma\nu=0$ (iv)
Also, considering the last four tilded spin coefficients we get:
$\epsilon-\bar{\epsilon}=0$ (139)
$\gamma-\bar{\gamma}=0$ (140)
$\alpha-\bar{\beta}=0$ (141)
The last three relations marks the difference compared to the simplifications made in the previous chapters, due to the exact same rotation. Based on the key relation of $\Psi_{1}=\kappa\mu-\sigma\pi$ and the relation $\alpha-\bar{\beta}=0$ we can easily derive the two following relations. The first one was obtained by the subtraction of the complex conjugate of NPE (d) with NPE (e) and the second one by the summation of the complex conjugate of NPE (r) with NPE (o).
$\pi\rho-\kappa\lambda=0$ (142)
$\Psi_{3}=\nu\rho-\lambda(\tau+\beta)$ (143)
By consideration of relations (124), (125) we choose to annihilate the spin coefficients $\kappa,\nu$ avoiding to correlate the elements of Killing tensor between themselves obtaining either $\lambda_{1}+\lambda_{2}+\lambda_{7}=0$ or $\lambda_{1}+\lambda_{2}-\lambda_{7}=0$. We made this choice since our goal for this chapter is to study a Killing tensor with one double eigenvalue. Moreover, the $K^{4}_{\mu\nu}$ form has 3 eigenvalues where $\lambda_{1}$ to be double and the others to be equal to $-(\lambda_{2}+\lambda_{7})$ and $-(\lambda_{2}-\lambda_{7})$ [^6] However, we briefly examined the case where $Q+1=0$ for $K^{4}_{\mu\nu}$ form proving that this special case admit Petrov types III, N, O in contrast with $K^{0}_{\mu\nu}$ of Chapter 5. The proof of the latter is not given in this paper.. The simplifications of spin coefficients till now are presented.
$\kappa=\nu=\epsilon-\bar{\epsilon}=\gamma-\bar{\gamma}=\alpha-\bar{\beta}=0$ (144)
Regarding this, three classes of solution emerge by relation (142). Either $\pi=0$ or $\rho=0$ or both.
8.2.1 Class I: $\rho\neq 0=\pi$
[…]
$\Psi_{2}-\Lambda=0$ (i)
$\Psi_{1}=0$ (ii)
$\Psi_{2}-\Lambda=0=\mu\rho-\sigma\lambda$ (iii)
$\mu\tau=0$ (iv)
where the last of them yields three possibilities. The first case of these possibilities is obtained when $\mu=0\neq\tau$.
$\lambda=0$ (n)
$\Psi_{2}=\Lambda=0$ (h)
$\Psi_{4}=0$ (j)
[…]
$\kappa=\nu=\pi=\alpha=\beta=\mu=\lambda=\epsilon-\bar{\epsilon}=\gamma-\bar{\gamma}=0$
$\Psi_{1}=\Psi_{2}=\Psi_{3}=\Psi_{4}=0$
The second case where $\mu=0=\tau$ is also of type N indicating that it is a subset of the first case.
We shall proceed to the third case which is the most interesting one since it is proved to be type D when $\mu\neq 0=\tau$. It is also important to show that the type D character of this solution emerged by combining the key relation (iii) $\mu\rho-\sigma\lambda=0$ with the BI (II) and (VII).
$0=3\Psi_{2}\rho-\lambda\Psi_{0}$ (II)
$0=\sigma\Psi_{4}-3\Psi_{2}\mu$ (VII)
Hence, dividing by parts we obtain the already known combination of the Weyl components as in the previous chapter.
$\Psi_{0}\Psi_{4}-(3\Psi_{2})^{2}=0$ (145)
The following relations encapsulate the nullified spins of the third case.
$\kappa=\nu=\pi=\tau=\alpha=\beta=\epsilon-\bar{\epsilon}=\gamma-\bar{\gamma}=0$
$\Psi_{1}=\Psi_{3}=0$
$\Psi_{0}\Psi_{4}=(3\Psi_{2})^{2}=(3\Lambda)^{2}$
The Table 5 contains the three solutions of Class I that we presented in this section.
$\kappa=\nu=\pi=\alpha=\beta=\epsilon-\bar{\epsilon}=\gamma-\bar{\gamma}=0$ (146)
It is neccessary to denote that for any of the following cases the previous spins annihilation hold.
Table 5: Class I: $\rho\neq 0=\pi$
Type N Type N Type D
$\mu=0\neq\tau$ $\mu=0=\tau$ $\mu\neq 0=\tau$
$\mu=\lambda=0$ $\mu=\lambda=\tau=0$ $\tau=0$
$\Psi_{0}\neq 0$ $\Psi_{0}\neq 0$ $\Psi_{0}\Psi_{4}-(3\Psi_{2})^{2}=0$
8.2.2 Class II: $\rho=0=\pi$
Initiating by $\kappa=\nu=\pi=\rho=0$ the key relation reads
$\Psi_{2}-\Lambda=0$ (i)
$\Psi_{1}=0$ (ii)
$\Psi_{2}-\Lambda=0=\sigma\lambda$ (iii)
$\mu\tau=0$ (iv)
the only non-conformally flat solution arises when $\mu\neq 0=\tau$. The NPE reads
$\sigma=0$ (a)
$\Psi_{0}=0$ (b)
$\Psi_{1}=0$ (c)
$\Psi_{3}=-\lambda\beta=0$ (i)
$\Psi_{2}=\Lambda=0$ (q)
$\epsilon(\mu-\bar{\mu})=0$ (l)
$-\Delta\lambda=\lambda(\mu+\bar{\mu}+2\gamma)+\Psi_{4}$ (j)
A type N solution is achieved only when $\alpha=\beta=0$ according to the NPE (i). All other scenarios result in type O solution, which does not pique our interest.
8.2.3 Class III: $\rho=0\neq\pi$
The key equation for this class is slightly different.
$\Psi_{2}-\Lambda=-\pi\tau$ (i)
$\Psi_{1}=-\pi\sigma$ (ii)
$\Psi_{2}-\Lambda=-\sigma\lambda$ (iii)
$\mu\tau=0$ (iv)
Implying the annihilation of $\sigma$ due to NPE (b) we get $\Psi_{0}=0$, also the key relation (i) reads
$\Psi_{2}-\Lambda=0=\pi\tau\hskip 5.69046pt\longrightarrow\hskip 2.84544pt\tau=0\hskip 2.84544pt;\hskip 4.26773pt\pi\neq 0$ (i)
Furthermore, implying the latter to NPE (k) and to NPE (q) we obtain $\Psi_{1}=0$ and $\Psi_{2}=\Lambda=0$ accordingly. Thus, the only non-zero Weyl component is $\Psi_{3}=-\lambda\beta$ indicating that this solution is of Petrov type III.
Theorem 5.
Petrov type N and D solutions with a geodesic, shearing and diverging null congruence admit $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$.
Theorem 6.
Petrov type N and III solutions with a geodesic, non-shearing and non-diverging null congruence admit $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$.
9 Discussion and Conclusions
The Study of the Canonical forms of Killing tensor is initiated by considering the possibility to obtain more interesting spacetimes using the canonical forms as an initial premise. In order to prove this statement properly we should rather have extracted the corresponding solutions that admit the canonical forms of Killing tensor in electro-vacuum, where the Einstein-Maxwell tensor $F_{\mu\nu}$ would be present and comparing to those of the work of Hauser-Malhiot [Hauser and Malhiot, 1976]. This work will be conducted in the immediate future. However, quite useful conclusions were extracted during this investigation (Vacuum with $\Lambda$) concerning this hypothesis.
In an attempt to address the aforementioned hypothesis and provide a proper answer, we sought to establish a rule to counterbalance the arbitrariness of the tetrad frame. In pursuit of solutions of the most general Petrov type, the authors believe that the transformation process must be as general as possible providing the most general Petrov type. As far as we know, the transformations applied do not follow any particular rule, and there is no any directive regarding the most general transformative process. For this reason, this study was operated scoping to imply only a null rotation around $l^{\mu}$, [^7] The same key relations would be obtained if $n^{\mu}$ was fixed. thus, the resolution process was attained under this rule. Regarding this, only $p=c+id$ and $a$ (boost) or $ib$ (twist) were annihilated due to the conservation of Killing tensor. Afterwards, none simplification takes place leaving the remained rotation parameters undetermined.
According to Remark 3, the null rotation is proved fruitful only when the reduced forms of Killing tensor are present since the invariant character of the Killing tensor would annihilate both $t=a+ib$ and $p=c+id$ “draining” the remaining tetrad freedom. Strictly speaking, the capitalization of the rotation parameters manages to gain the key relations which basically reveal the classes of our solutions. Differently, we could take advantage of the free rotation parameter $t$ correlating the spin coefficients within themselves (Lorentz transformations) only. In other words, the only way to obtain simplifications is either to employ reduced Killing tensor or to imply Lorentz transformations correlating the spin coefficients between themselves which is equivalent to the determination of $t$.
The latter choice falls into the concept of the symmetric null tetrads that was manifested in [Debever and McLenaghan, 1981]. Using the free rotation parameters $a$ and $b$ Debever managed to obtain the following simplifications that led him to the most general type D solution with a geodesic, shear-free null congruence [Debever et al., 1984] assuming that the PND of Einstein-Maxwell tensor aligned to the PND of Weyl tensor. Namely, the following relations determine the rotation parameters $a$ and $b$
$\pi=e\tau$
$\mu=e\rho$
$\gamma=e\epsilon$
The interesting part arises when we derived the Petrov types of the extracted solutions assuming the existence of the Killing tensor of our Paradigm both ways. Firstly we implied the symmetric null tetrad concept, afterwards the null rotation methodology of Chapter 3. With fundamental derivations it is really easy to prove that using the symmetric null tetrad concept one extracts a unique type D solution in vacuum with $\Lambda$. In fact, this type D solution is a limitation in vacuum with $\Lambda$ of Debever’s solution [Debever et al., 1984] and Papakostas’ solution [Papakostas, 1998], wherein both solutions were obtained based on the aforementioned concept.
Although, applying the methodology of Chapter 3 we result in two conformally flat solutions  [^8] The proof of this statement is not given here, although it is easy for the reader to verify this using the IC of the Paradigm (relations (11) in [Papakostas, 1998])., wherein the implication of the null rotation just annihilates the parameter $p$. Basically this outcome shows that we obtain more general Petrov types solutions (type D) abolishing the arbitrariness of the tetrad. Subsequently, a question arises, which is the most general transformation that we must apply in order to acquire the most general Petrov type solution?
Moving forward, another important matter was brought to surface by this study, is that the canonical forms appear to be more fruitful instead of the Killing tensor of our Paradigm satisfying the conjecture of Hauser-Malhiot [Hauser and Malhiot, 1976]. This is evident by the vast variety of the different Petrov types were obtained assuming the $K^{2},K^{3}$ Killing forms and $K^{4}$ form by extension  [^9] Once more, we denote that $K^{4}$ form is not a typical canonical form but it is originated by the reduction of $K^{1},K^{2},K^{3}$ where $\lambda_{0}=0$.. Also, in reference to the work of Van den Bergh [Van den Bergh, 2017] our type D solutions, one with a shear-free and non-geodesic null congruence (Theorem 4) and the other with a shearing and geodesic null congruence (Theorem 5) might be new type D solutions. In fact, they could be possible reductions (in vacuum with $\Lambda$) of the undiscovered solutions with PND of Einstein-Maxwell tensor aligned (or non-aligned) with the PND of Weyl tensor.
Finally, we hope that a similar study of the canonical forms in electro-vacuum with $\Lambda$ will be able to address most of the questions posed. In this future work, we also aim to investigate whether the generality of the canonical forms of the Killing tensor leads to more general canonical forms of Weyl tensor.